### v1

# Blessing of Multilinguality: A Systematic Analysis of Multilingual In-Context Learning

Abstract
While multilingual large language models generally perform adequately, and sometimes even rival English performance on high-resource languages (HRLs), they often significantly underperform on low-resource languages (LRLs; Costa-jussà et al., 2022). Among several prompting strategies aiming at bridging the gap, multilingual in-context learning (ICL; Shi et al., 2023) has been particularly effective when demonstration in target languages is unavailable. However, there lacks a systematic understanding when and why it works well.
In this work, we systematically analyze multilingual ICL, using demonstrations in HRLs to enhance cross-lingual transfer. We show that demonstrations in mixed HRLs consistently outperform English-only ones across the board, particularly for tasks written in LRLs. Surprisingly, our ablation study show that the presence of irrelevant non-English sentences in the prompt yields measurable gains, suggesting the effectiveness of multilingual exposure itself. Our results highlight the potential of strategically leveraging multilingual resources to bridge the performance gap for underrepresented languages.
Blessing of Multilinguality: A Systematic Analysis of Multilingual In-Context Learning
Yilei Tu Andrew Xue Freda Shi David R. Cheriton School of Computer Science, University of Waterloo 200 University Ave W, Waterloo, ON N2L 5Z5, Canada {yileitu,a8xue,fhs}@uwaterloo.ca
1 Introduction
In-context learning (ICL; Brown et al., 2020) has become a widely adopted technique in natural language processing with large language models (LLMs; Touvron et al., 2023; Chowdhery et al., 2023; Dubey et al., 2024; Yang et al., 2024a, b, inter alia), which enables LLMs to learn to solve problems by analogy from a few input-output examples (i.e., demonstrations) without updating model parameters. As a generic method, ICL has also been effective in improving the cross-lingual performance of multilingual LLMs (MLLMs; Winata et al., 2021; Ahuja et al., 2023; Shi et al., 2023; Razumovskaia et al., 2024; Asai et al., 2024).
Prior work has introduced two common ICL strategies for non-English languages: (i) translating the target question into English and performing English-only ICL (Shi et al., 2023; Ahuja et al., 2023; Razumovskaia et al., 2024, inter alia), and (ii) providing demonstrations in the target language (in-language demonstrations; Fu et al., 2022; Qin et al., 2023; Huang et al., 2023; Zhang et al., 2024b). Both strategies have critical shortcomings: (i) may suffer from information loss in translation due to nuanced language gap (Zhang et al., 2023; Poelman and de Lhoneux, 2024) or the unavailability of high-quality translation systems for extremely low-resource languages (LRLs), whereas (ii) may become infeasible due to data scarcity in LRLs. As alternatives, when presenting LLMs with the problems in the target language, English demonstrations (Fig. 1(a)) lead to poor performance on LRLs, whereas demonstrations in multiple high-resource languages (HRLs; Fig. 1(b)) can be more effective, even when there are little alphabetical overlap between the demonstration and target languages (Shi et al., 2023); however, the underlying reasons on why it works remain unclear.
In this work, we systematically analyze multilingual ICL through a set of controlled experiments. Each test question is paired with a set of semantically equivalent demonstrations, while the demonstration languages vary according to different ICL modes (Tab. 1). We compare the performance differences across four ICL modes (§§ 4.1 and 4.2): English, individual-HRL, multilingual (i.e., mixed HRLs), and in-language demonstrations (Fig. 2). To disentangle the impact of demonstration language from other confounding factors (e.g., interactions between demonstration languages and in-domain demonstrations), we conduct additional control experiments by adding irrelevant sentences in various languages into English-only in-domain demonstrations (§ 4.3). We find that
In-context demonstrations in HRLs, especially in languages with non-Latin writing system such as Chinese and Japanese, can more effectively transfer knowledge compared to English ICL mode, leading to performance improvement on answering questions in all languages, especially in LRLs. This finding is generalizable enough across different domains and various LLMs.
Demonstrations in mixed HRLs is more robust and effective compared to that in a single HRL, in terms of average accuracy boosting on different tasks. This strategy is favored.
Surprisingly, simply introducing another non-English language (not necessarily in the demonstration) in the prompt could lead to performance improvement, albeit the improvement is not as significant as the aforementioned strategies.
2 Background: Multilingual ICL
In this section, we review the basic approaches of ICL with instruction-tuned LLMs (§ 2.1) and introduce the multilingual prompting modes (§ 2.2) that we evaluate in this work.
2.1 ICL for Instruction-Tuned LLM
Instruction-tuned LLMs (Ouyang et al., 2022; Mishra et al., 2022; Wang et al., 2022; Wei et al., 2022a) generally possess the capability to follow task instructions (i.e., system prompt), which are usually coupled with ICL to fully elicit their capability (Wei et al., 2022b, inter alia). Formally, denote a training set by and a test set in the same domain, where is a task question (as model input). An ICL prompt for a test question has three core components: (1) a system prompt that describes the task and specifies the expected output format, (2) sample input-output pairs (-shot) from the training set that provide in-context demonstrations, and (3) a verbalizer mapping each ground truth label to a textual representation, which may also include reasoning steps (i.e., chains of thoughts, or CoT in short; Wei et al., 2022b). In summary, an ICL prompt for can be written as:
where is the string concatenation operator with a special end-of-turn token as the delimiter. The LLM with parameters , denoted as , then generates the response given : .
2.2 Multilingual Prompting Modes
We extend our notations as follows to adapt to the multilingual settings. A training set with languages is denoted by , where the split consists of examples for any . The same applies to the test dataset , with each language-specific split consisting of examples. Without further specification, we assume that the training examples at the same index are semantically equivalent across languages.
The ground-truth labels in quantitative LLM benchmarks (Ponti et al., 2020; Cobbe et al., 2021, inter alia) are typically language-agnostic (such as numbers) or represented in a single word (such as Yes/No). In such cases, the verbalizer is an identity. For answers requiring reasoning steps, is CoT in English, as there has been strong evidence that MLLMs performs better when generating English (Shi et al., 2023; Qin et al., 2023; Huang et al., 2023, inter alia). For the same reasons, is always presented in English as well.
Following Ahuja et al. (2023) and Shi et al. (2023), we evaluate MLLMs via several different prompting strategies (ICL modes) in this work:
The English mode. The demonstrations are always in English (Figs. 1(a) and 2a).
The Monolingual mode(s). The demonstrations are always presented in a single non-English high-resource language such as Chinese (Fig. 2b).
The Multilingual mode. From a predefined list of high-resources languages , languages are randomly selected, which, together with the sampled indices, determine the contents and languages of the demonstrations (Figs. 1(b) and 2c).
The Native mode. The demonstrations are in the same language as the test question (Fig. 2d).
3 Experiment Setups
Models. We evaluate state-of-the-art instruction-tuned LLMs with about 8 billion parameters, which have officially claimed multilingual capabilities in model release: Llama3-8B-Instruct, Llama3.1-8B-Instruct (Dubey et al., 2024), Qwen2-7B-Instruct (Yang et al., 2024a), Qwen2.5-7B-Instruct (Yang et al., 2024b), Mistral-NeMo-12B-Instruct (MistralAI, 2024) and Aya-Expanse-8B (Dang et al., 2024). For additional references, we evaluate OpenAI closed-sourced commercial models, including GPT3.5-turbo (OpenAI, 2022) and GPT4o-mini (OpenAI, 2024). Detailed model cards can be found in § A.1.
| Dataset | Domain | Expected | #Languages | Language Split Size | En Avg Word | Parallel |
| Datapoint Example | Output | (#HRL #LRL) | — Training/Test | Count | ||
[…]
| See Fig. 1 for examples. | ||||||
[…]
| Premise: The man turned on the faucet. What happened as a RESULT? | ||||||
| Hypothesis 1: The toilet filled with water. Hypothesis 2: Water flowed from the spout. Answer: 2. | ||||||
[…]
| Sentence 1: What did you *get* at the toy store? Sentence 2: She didn’t *get* his name when they met the first time. | ||||||
| Question: Is the word “get” (marked with *) used in the same way in both sentences above? Answer: No. | ||||||
Datasets. We evaluate the models using multilingual benchmarks from three distinct domains: (1) MGSM Shi et al. (2023), a benchmark of 250 grade-school math problems sampled from the English GSM8K (Cobbe et al., 2021) and translated into 10 additional languages by expert native speakers. (2) XCOPA Ponti et al. (2020), a commonsense reasoning benchmark that extends the COPA dataset (Roemmele et al., 2011) to additional languages.111In this work, we merge the English COPA and XCOPA datasets together, which we still refer to as XCOPA. (3) XL-WiC Raganato et al. (2020), a cross-lingual word-in-context understanding dataset spanning languages, where models are expected to tell whether a polysemous word retains the same meaning in two contexts.
MGSM and XCOPA are parallel where each corresponding datapoint across different language splits contains semantically equivalent content, allowing us to minimize semantic confounders in our experimental design. XL-WiC is language-specific and translation-variant thus naturally non-parallel. Dataset properties and examples are in Tabs. 1 and 6. Details of data curation are in § A.2.
Languages. Languages with large-scale digitized data resources on web are known as high-resource languages (HRLs; Bender, 2019), which are exemplified by English, Spanish, Chinese, among others. In contrast, low-resource languages (LRL) have scarce accessible data Costa-jussà et al. (2022). However, a universal standard for dichotomizing languages as either high- or low-resource has not been set (Bender, 2019; Joshi et al., 2020; Hedderich et al., 2021). Moreover, none of the models we evaluate has disclosed the language distributions in their training corpora. As a workaround, we define our preset HRL list as the union of the 20 most frequent languages in Llama2 Touvron et al. (2023) and PaLM Chowdhery et al. (2023), and classify languages out of the HRL list as LRLs. Details of preset HRL list can be found in § A.3.
Prompts. Following Shi et al. (2023), we use examples for demonstration for any test question. For each multilingual dataset (Tab. 1), we first sample index lists of length all at once, where the index range is We allocate the -th index list to the set of test questions for the same index across all language splits. The training-set index list, together with the specified languages,222 For the Multilingual mode, we apply the same sampling procedure to generate HRL code lists of length , drawing from the available HRLs in the dataset. jointly determines the content and language of the demonstration for each testing example (Fig. 2). This approach both ensures linguistic diversity for multilingual prompting and, whenever applicable, mitigates confounding factors that come with semantic inconsistency across examples. All interface words (see Tab. 1) are in the same language as the examples rather than in English.
Inference and metrics. Throughout this work, we use greedy decoding for inference, selecting the token with the highest probability at each step. For MGSM in need of CoT, we set the maximum token length to 500; for XCOPA and XL-WiC, we set it to 10, as we expect the answers to be short. We use exact match accuracy as our evaluation metric: for MGSM, we extract the last numeral in the response. For XCOPA and XL-WiC, we extract label (expected output) in the response (Tab. 1).
4 ICL Mode Evaluation
4.1 Multilingual Prompts Surpass English
3: Llama3-8B-Instruct;
3.1: Llama3.1-8B-Instruct;
2: Qwen2-7B-Instruct;
2.5: Qwen2.5-7B-Instruct;
: Mistral-NeMo-12B-Instruct;
: Aya-Expanse-8b;
3.5: GPT3.5-turbo;
4om: GPT4o-mini. | Acc | MGSM | XL-WiC | XCOPA | ||||||||
| bn | sw | te | th | LRL Avg | bg | et | fa | hr | LRL Avg | LRL Avg | |
| English | 52.40 | 67.20 | 39.60 | 69.20 | 57.10 | 51.54 | 44.62 | 29.74 | 51.79 | 44.42 | 55.91 |
| Multilingual | 68.00 | 68.80 | 55.60 | 71.60 | 57.69 | 55.13 | 57.95 | 56.92 | |||
| Native | 67.20 | 72.40 | 58.00 | 76.40 | 61.79 | 62.05 | 62.31 | 57.18 | |||
| English | 57.20 | 20.80 | 23.20 | 73.60 | 43.70 | 28.21 | 56.92 | 63.33 | 54.87 | 48.46 | 62.29 |
| Multilingual | 64.80 | 28.40 | 23.60 | 73.20 | 53.59 | 58.46 | 60.26 | 54.36 | |||
| Native | 72.00 | 32.40 | 40.40 | 76.80 | 55.13 | 59.23 | 63.08 | 54.62 | |||
| English | 39.60 | 63.60 | 12.80 | 60.80 | 44.20 | 54.36 | 54.62 | 54.10 | 51.79 | 53.72 | 63.43 |
| Multilingual | 54.40 | 68.00 | 24.40 | 57.20 | 52.82 | 60.00 | 54.36 | 56.41 | |||
| Native | 57.20 | 73.60 | 30.00 | 58.80 | 54.62 | 59.49 | 58.46 | 60.26 | |||
Results. We first compare the -shot performance with English, Multilingual and Native ICL modes on open-sourced MLLMs and commercial OpenAI models (Fig. 3), with Tab. 2 presenting the detailed performance of three selected MLLMs across various LRLs. Overall, Multilingual mode outperforms English mode, both for individual LRLs and on average. In out of cases (Fig. 3), Multilingual mode achieves higher accuracy than English one. This phenomenon is evident even for GPT4o-mini, one of the currently strongest LLMs Chiang et al. (2024), and for HRLs as well (see Tabs. 11, 12 and 13 in § B.1 for HRL accuracies). Extending the results of Shi et al. (2023) that Multilingual mode generally outperforms English mode for PaLM (Chowdhery et al., 2023) and Codex (Chen et al., 2021), our results confirm this trend is a general phenomenon across various MLLMs and datasets.
Outstanding performance of the Native mode and the practical unfeasibility. Admittedly, in out of comparisons, Native mode outperforms Multilingual mode (Fig. 3), which aligns with the machine-learning intuition that in-domain data, in terms of both genre and language, are more promising for model performance (Liu et al., 2024). However, domain-specific datasets for LRLs are often difficult to obtain due to the scarcity of native speakers or professional translators (Costa-jussà et al., 2022; NLLB-Team, 2024); therefore, in practice, it is usually challenging to provide high-quality demonstrations in the same language and domain as the test question. In contrast, annotations makes the HRL-Multilingual mode more feasible in many scenarios.
Hypothesis Tests. To verify whether the improvement is statistically significant, we conduct McNemar’s test McNemar (1947)—the null hypothesis means no significant accuracy difference between the baseline (English) and the compared mode. Let denote the number of cases where the baseline is correct while the compared mode is incorrect, denotte the number of cases where the baseline mode is incorrect while the compared mode is correct. We calculate the corrected version Edwards (1948) of the McNemar’s statistic:
which has a chi-squared distribution with one degree of freedom. Significant -test results provide strong evidence to reject the null hypothesis of no accuracy improvement. Our results results (Tab. 2 and Tabs. 14, 15 and 16 in § B.1) indicate that both Multilingual and Native modes significantly outperform the English mode.
4.2 Ablation Study: Non-English Monolingual Prompts Are Effective
With the success of the Multilingual mode, we investigate whether the improvement comes from the introduction of multiple languages or simply from a single non-English HRL. Specifically, we compare several Monolingual modes including Chinese (for all three datasets), French, Japanese (both for MGSM and XL-WiC) and Italian (for XCOPA). We select French and Italian because they are both European languages and thus share considerable subword overlap with English; in contrast, Chinese and Japanese exhibit little subword overlap with Latin-script languages, but there is a substantial overlap between them two due to their shared use of Chinese characters (or kanji). This language selection allows us to analyze simultaneously the impact of the writing system (or subword overlap) on ICL performance.
Results. All HRL-Monolingual modes outperform English in a considerable number of comparisons (Fig. 4). This finding also holds for HRL evaluations (see Tabs. 11, 12 and 13 in § B.1 for raw accuracies of each language). Among all Monolingual modes, Chinese performs the best, matching Multilingual mode with 17 out of 21 comparisons when accuracy surpasses that of English. Japanese also frequently outperforms English. Extending the findings of Turc et al. (2021) that non-English languages are more effective than English in pretraining and fine-tuning based cross-lingual transfer, our results suggest that non-English languages, particularly those with non-Latin scripts, may be more effective under the prompting scheme as well.
However, the Multilingual mode exhibits stronger robustness, outperforming Chinese in 14 out of 21 LRL cases. Same trends applies to HRLs. The results of the hypothesis test further confirm the robustness: for both LRL and HRL splits, the Multilingual mode exhibits the highest number of significant results relative to other Monolingual modes (Tabs. 14, 15 and 16 in § B.1). Intuitively, we hypothesize that Multilingual mode functions like an “average” of individual HRL-Monolingual modes, making it most robust, and thus it outperforms English most frequently and achieves the highest overall average accuracy.
Following Tang et al. (2024), we further identify ICL-mode-specific neurons and find that the neurons activated by Multilingual overlap most with those activated by Native among other modes. This further explains why Multilingual could achieve performance comparable to Native. See App. C for details.
4.3 Ablation Study: Merely Introducing New Language(s) Enhances ICL Performance
English ICL Mode Multilingual Context-Irrelevant Sentences
(English CIS-Multi)
After showing that including non-English in the prompts can improve ICL performance, a natural follow-up question arises: does the gain come from the mere presence of non-English languages, or the interaction between the target language and the in-topic examples? To distinguish these two setups, we prepend a context-irrelevant sentence (CIS) before each ICL demonstration, which is unrelated to the current domain and can be in any language. Analogous to the ICL modes introduced in § 2.2, CIS resembles the construction of those modes with the same sampling strategy (§ 3). For example, based on the English mode, we could prepend a set of multilingual CIS, which augments Fig. 2a into Fig. 5. We denote such setting by “English CIS-Multi”. This naming convention applies to other settings accordingly.
We use sentences from FLORES-101 Goyal et al. (2022) as the source of our CIS, which provides parallel Wikipedia sentences in multiple languages, a fairly distant genre to all evaluation datasets. We filter sentence-parallel sets with English word counts ranging from 10 to 15 as our sampling pool , a small range compared to target datasets (Tab. 1), to mitigate the risk of introducing too much noise. Details of FLORES-101 can be found at § A.2.
Results and analysis. We perform hypothesis testing as in Tab. 2 to compare different CIS settings with English CIS-En (Tab. 3)—English CIS-En exhibit a small drop compared to English only, indicating that our filtration effectively controls the negative impact of noise to a tolerable level. We find that introducing a single non-English language generally improves ICL performance in most cases ; however, the improvement is only statistically significant in cases. This reveals that simply introducing a language can lead to a modest improvement in MLLM performance, but it is more pronounced when in-topic demonstrations in another language (Monolingual modes) are incorporated. More hypothesis test results for both LRL and HRL splits can be found in Tabs. 20, 21 and 22 in § B.2.
Introducing multiple languages (CIS-Multi, Fig. 5) is slightly more promising than CIS in a single language, with of cases showing improvement, of which are significant. We further conduct experiments by prepending multilingual CIS to the Multilingual mode (Multilingual CIS-Multi, Tab. 3). The performance of Multilingual CIS-Multi is not significantly lower than that of English CIS-Multi; in more than half cases, it significantly outperforms English CIS-Multi. This concludes that the significant improvement from English mode to Multilingual mode is attributed to the incorporation of multiple languages with in-topic demonstrations.
| LRL AvgΔ↑↓ | MGSM | XL-WiC | XCOPA |
| English | 64.20 | 57.37 | 46.23 |
| English CIS-En | 61.00 | 56.73 | 48.37 |
| English CIS-Fr | |||
| English CIS-Ja | |||
| English CIS-Zh | |||
| English CIS-Multi | |||
| Multilingual CIS-Multi | |||
| English | 57.10 | 44.42 | 55.91 |
| English CIS-En | 55.90 | 47.88 | 55.46 |
| English CIS-Fr | |||
| English CIS-Ja | |||
| English CIS-Zh | |||
| English CIS-Multi | |||
| Multilingual CIS-Multi | |||
| English | 43.70 | 48.46 | 62.29 |
| English CIS-En | 43.00 | 50.90 | 62.31 |
| English CIS-Fr | |||
| English CIS-Ja | |||
| English CIS-Zh | |||
| English CIS-Multi | |||
| Multilingual CIS-Multi | |||
| English | 59.40 | 57.82 | 64.69 |
| English CIS-En | 59.40 | 59.04 | 64.20 |
| English CIS-Fr | |||
| English CIS-Ja | |||
| English CIS-Zh | |||
| English CIS-Multi | |||
| Multilingual CIS-Multi | |||
| English | 57.00 | 51.54 | 62.26 |
| English CIS-En | 60.70 | 49.62 | 61.00 |
| English CIS-Fr | |||
| English CIS-Ja | |||
| English CIS-Zh | |||
| English CIS-Multi | |||
| Multilingual CIS-Multi | |||
| English | 29.40 | 58.78 | 26.54 |
| English CIS-En | 27.80 | 57.82 | 23.20 |
| English CIS-Fr | |||
| English CIS-Ja | |||
| English CIS-Zh | |||
| English CIS-Multi | |||
| Multilingual CIS-Multi | |||
4.4 Translation-Based Performance
Mirroring the translation-training (Hu et al., 2020, inter alia) setup, existing work suggest that translating from LRLs into English and prompting with the translation results may yield better results (Ahuja et al., 2023, inter alia). While translation is not the main focus of our work, we conduct experiment to compare the performance of translation-based strategies for reference.
Strategies. We test two translation strategies for baseline comparison. (1) Transl-LangEn: Translating test questions in other languages in the English ICL mode (Fig. 2a) into English. (2) Transl-EnLang: Translating demonstrations in English ICL mode (Fig. 2a) into the source language of the current test question, which mirrors the Native mode (Fig. 2d). We use the Google Cloud Translation API for translation.333https://cloud.google.com/translate/
| Avg Acc | MGSM | XCOPA | ||
| LRL | HRL | LRL | HRL | |
| Multilingual | 66.00 | 79.20 | 66.11 | 89.64 |
| Native | 68.50 | 79.43 | 71.63 | 90.80 |
| Transl-LangEn | 60.00 | 74.80 | 76.74 | 89.68 |
| Transl-EnLang | 68.60 | 80.63 | 70.49 | 90.04 |
| Multilingual | 59.50 | 86.97 | 64.63 | 91.84 |
| Native | 60.30 | 87.31 | 67.69 | 92.64 |
| Transl-LangEn | 63.40 | 78.63 | 80.49 | 91.52 |
| Transl-EnLang | 59.90 | 86.97 | 66.23 | 92.48 |
| Multilingual | 63.30 | 81.09 | 62.94 | 88.16 |
| Native | 65.80 | 81.26 | 70.91 | 90.28 |
| Transl-LangEn | 61.60 | 76.00 | 77.83 | 89.32 |
| Transl-EnLang | 66.60 | 82.00 | 68.46 | 90.52 |
Analysis. For LRLs, Transl-LangEn sometimes outperform the Native mode, while Transl-EnLang performs comparably to the Multilingual mode, but falls short of the Native mode performance (Tab. 4). These results resonate with the phenomena that the translation quality of LRLEn is generally higher than that of the reversed direction Fan et al. (2021); Goyal et al. (2022); Costa-jussà et al. (2022).
For HRLs, however, Transl-LangEn underperforms to Native and, sometimes, even Multilingual (e.g., on MGSM), suggestting that if a language is sufficiently well-trained, generating responses directly in that language is more effective than translating into English before inference. These two findings highlight that MLLMs approach the ideal of being equally capable in HRLs Liu et al. (2024), but are still undertrained on LRLs, making translation into English a favored strategy.
In agreement with Poelman and de Lhoneux (2024), we would like to note that even if the task performance of translation-based strategies are the best, the ultimate goal of multilingualism is not just about optimizing task-specific performances. A universal language model should be able to understand and generate text in all languages, instead of relying on specific language(s) as an intermediary. On the other hand, due to the loss of semantic nuances, grammatical structures and cultural context, translation-based strategies may not be the best choice for tasks heavily reliant on language-specific nuances Liu et al. (2024).
5 Related Work
Prompt engineering. Instruction tuning aligns LLMs more closely with human instructions Ouyang et al. (2022); Mishra et al. (2022); Wei et al. (2022a); Askell et al. (2021); Wang et al. (2022, 2023). Concurrently, numerous prompting strategies have been developed Liu et al. (2023) and shown to consistently enhance the performance of instruction-tuned LLMs, such as in-context learning Brown et al. (2020); Min et al. (2022) and chain-of-thought Wei et al. (2022b); Kojima et al. (2022). These prompting strategies are proven effective in multilingual tasks as well Winata et al. (2021); Lin et al. (2022); Shi et al. (2023).
Multilingual ICL. For languages of templates, demonstration and sample question in native language is conventionally inserted into a predefined English template Lin et al. (2022); Fu et al. (2022); Qin et al. (2023); Huang et al. (2023); Ahuja et al. (2023); Zhang et al. (2024b). Poelman and de Lhoneux (2024) critiques this widespread misuse of English as the interface language. Qin et al. (2023); Huang et al. (2023); Zhang et al. (2024b) guide models to “think” and generate CoT in English, regardless of the input language, leading to improved performance for generation tasks compared to “thinking” in other language(s). Sclar et al. (2024); Zhang et al. (2024a) highlight that models are sensitive to those templates. For languages of demonstrations and test questions, Shi et al. (2023); Ahuja et al. (2023) conclude that in-language demonstrations outperform English demonstrations. Etxaniz et al. (2024); Liu et al. (2024) suggest translating questions from LRLs into English can improve performance.
6 Conclusion and Discussion
This work systematically analyzes multiple ICL strategies for MLLMs, and confirms that the presence of multiple languages is an effective strategy across multiple MLLMs. This improvement is partially due to the inclusion of non-English languages in the prompting, and partially due to the in-topic demonstrations in non-English languages, which together strengthen the models’ cross-lingual transfer capabilities, particularly the capability to process LRLs. Our work echoes with Turc et al. (2021)—who suggest that HRLs other than English excel in the pretraining-finetuning framework—in the in-context learning framework, highlighting the importance of language inclusivity.
We are in agreement with Costa-jussà et al. (2022) and Liu et al. (2024) that an ideal language-universal LLM should be equally capable in all languages. Beyond this belief, we found that non-English languages may better elicit the potential of MLLMs. Although these observations remain in using HRLs for LRL processing, our results strongly support the call for greater research investment in enhancing MLLM capabilities for a broader range of languages.
Limitations
This paper treats multilingual LLMs as black-box models, drawing the findings and conclusions based solely on their input-output behavior. Hence we have not interpreted the internal mechanism how multilinguality could affect MLLM’s “thinking” process and its manifested performance. We have briefly touched the impact of demonstrations in different languages on the MLLM performance. However, we do not conduct a thorough empirical analysis to identify which specific linguistic characteristics (e.g., writing systems, grammatical structures, or linguistic relatedness) contribute to the observed performance differences.
Reproducibility
Our code is available at https://github.com/yileitu/multilingual_icl.
Appendix A Experiment Setup
A.1 Model
| Model Name | Scale | Instruct? | Open-Source? | Checkpoint |
[…]
All model checkpoints we use and their properties are listed in Tab. 5.
A.2 Dataset
| Dataset | HRL | LRL | Source |
| MGSM | de, en, es, fr, ja, ru, zh | bn, sw, te, th | juletxara/mgsm |
| XCOPA | en, id, it, tr, zh | et, ht, qu, sw, ta, th, vi | English COPA & cambridgeltl/xcopa |
| XL-WiC | da, de, en, fr, it, ja, ko, nl, zh | bg, et, fa, hr | pilehvar.github.io/xlwic/ |
After preprocessing, datapoints for each language split are stored in a single JSON file. Tab. 6 summarizes the supported languages for each dataset and the sources from which they are obtained.
MGSM
The original dataset consists of parallel datapoints across all language splits and training/test splits of the same size. Example datapoint and the Chat Template for few-shot demonstrations can be found in Fig. 6.
XCOPA
The datapoints in the training split of XCOPA are parallel to the last 100 datapoints in the English COPA development split. Therefore, we exclude the first datapoints from the development split of English COPA. The test splits of both datasets are parallel and contain the same number of datapoints. We then merge both into our XCOPA dataset. In XCOPA, there are two types of questions: “cause” and “effect”, each corresponding to a distinct template, as shown in Figs. 7(a) and 7(b).
XL-WiC
This benchmark is designed to determine whether a specific word in a given language has the same meaning in two different sentences. As a result, the dataset is inherently non-parallel. Among all language splits, Estonian (et) contains the fewest datapoints, with in the training split and in the test split. For all other languages, we randomly subsample to match the size of the Estonian split to satisfy the demonstration sampling requirements outlined in Fig. 2. To leverage the attention mechanism of transformers Vaswani et al. (2017), we add asterisks around the target word in both sentences to indicate that the LLM needs to disambiguate the meaning of that specific word. An example datapoint is shown in Fig. 8.
COMBINED
Identifying specific neurons depends on the nature of the input corpus. Since language is inherently conjugate with the task, and our focus is on language-specific neurons rather than task-specific neurons, it is necessary to input all three datasets into the LLM to eliminate the confounding factor of the task domain. To balance the three datasets, we subsample their test splits to only datapoints each. To balance the number of language splits across datasets, we excluded two HRL splits, Korean (ko) and Dutch (nl), from the XL-WiC dataset. This ensures that all three (sub-)datasets have languages for combination. Additionally, for MGSM, we limit the answers to only include the final numeric result without the CoT reasoning. This approach ensures that the total number of datapoints and the overall token count are roughly the same across the original three datasets.
FLORES-101
This machine translation benchmark comprises sentences extracted from English Wikipedia, spanning diverse topics and domains. These sentences were translated into languages by professional translators via a carefully controlled process. The Wikipedia domain is largely unrelated to the domains of the three datasets we evaluate (math, commonsense reasoning, and word disambiguation). Therefore, we choose FLORES-101 as our source pool of irrelevant sentences. Note that in this benchmark, each datapoint carries the same semantic meaning across all language splits. We do not want irrelevant sentences to affect the LLM’s understanding of the original task excessively, thus we select sentences with word counts between and in the English split, introducing limited noise. CIS are drawn form the filtered FLORES-101 dataset. An example is provided in Fig. 9.
A.3 Language
A.3.1 High-Resource Language List
| Code | Language | Percentage |
| en | English | |
| de | German | |
| fr | French | |
| sv | Swedish | |
| zh | Chinese | |
| es | Spanish | |
| ru | Russian | |
| nl | Dutch | |
| it | Italian | |
| ja | Japanese | |
| pl | Polish | |
| pt | Portuguese | |
| vi | Vietnamese | |
| uk | Ukrainian | |
| ko | Korean | |
| ca | Catalan | |
| sr | Serbian | |
| id | Indonesian | |
| cs | Czech | |
| fi | Finnish |
| Code | Language | Tokens (B) | Percentage |
| en | English | 578.064 | |
| de | German | 25.954 | |
| fr | French | 24.094 | |
| es | Spanish | 15.654 | |
| pl | Polish | 10.764 | |
| it | Italian | 9.699 | |
| nl | Dutch | 7.690 | |
| sv | Swedish | 5.218 | |
| tr | Turkish | 4.855 | |
| pt | Portuguese | 4.701 | |
| ru | Russian | 3.932 | |
| fi | Finnish | 3.101 | |
| cs | Czech | 2.991 | |
| zh | Chinese | 2.977 | |
| ja | Japanese | 2.832 | |
| no | Norwegian | 2.695 | |
| ko | Korean | 1.444 | |
| da | Danish | 1.387 | |
| id | Indonesian | 1.175 | |
| ar | Arabic | 1.091 |
To the best of our knowledge, we find two multilingual LLMs—
Llama2 Touvron et al. (2023) and
PaLM Chowdhery et al. (2023)—that publicly report the language distribution used during pretraining. The top 20 languages and their percentages for each model are listed in Tab. 7 and Tab. 8, respectively. We take the union of these languages to form what we consider a high-resource language list (in terms of the richness in the LLM pretraining dataset), including the following languages:
The high overlap between the top languages of the two LLMs further supports the rationale for applying this list to other LLMs.
A.3.2 Languages We Evaluate
Tab. 9 presents the union of languages supported by the three datasets we evaluated (§ 3, datasets). Based on whether a language appears in the high-resource language list, we categorized the languages into HRL and LRL groups, with classified as HRL and as LRL. Among them, only English and Chinese are present in all three datasets.
It is worth highlighting that the LRLs span distinct writing systems and language families. This diversity suggests that when tokenizing inputs in these LRLs, they are unlikely to share common tokens with HRLs ( out of use the Latin script). Consequently, this limits the MLLMs’ ability to leverage shared subwords or similar syntax structures for cross-lingual transfer across LRLs.
| Code | Name in English | HRL/LRL | In Which Dataset(s) | Writing System | Family |
| bg | Bulgarian | Low | XL-WiC | Cyrillic | Indo-European |
| bn | Bengali | Low | MGSM | Bengali–Assamese | Indo-European |
| da | Danish | High | XL-WiC | Latin | Indo-European |
| de | German | High | MGSM, XL-WiC | Latin | Indo-European |
| en | English | High | MGSM, XL-WiC, XCOPA | Latin | Indo-European |
| es | Spanish | High | MGSM | Latin | Indo-European |
| et | Estonian | Low | XCOPA | Latin | Indo-European |
| fa | Persian | Low | XL-WiC | Arabic | Indo-European |
| fr | French | High | MGSM, XL-WiC | Latin | Indo-European |
| hr | Croatian | Low | XL-WiC | Latin | Indo-European |
| ht | Haitian | Low | XCOPA | Latin | French Creole |
| id | Indonesian | High | XCOPA | Latin | Austronesian |
| it | Italian | High | XL-WiC, XCOPA | Latin | Indo-European |
| ja | Japanese | High | MGSM, XL-WiC | Kana & Chinese Characters | Japonic |
| ko | Korean | High | XL-WiC | Hangul | Koreanic |
| nl | Dutch | High | XL-WiC | Latin | Indo-European |
| qu | Quechua | Low | XCOPA | Latin | Quechumaran |
| ru | Russian | High | MGSM | Cyrillic | Indo-European |
| sw | Swahili | Low | MGSM | Latin | Niger–Congo |
| ta | Tamil | Low | XCOPA | Tamil | Dravidian |
| te | Telegu | Low | MGSM | Telegu | Dravidian |
| th | Thai | Low | MGSM, XCOPA | Thai | Kra–Dai |
| tr | Turkish | High | XCOPA | Latin | Turkic |
| vi | Vietnamese | Low | XCOPA | Latin | Austroasiatic |
| zh | Chinese | High | MGSM, XL-WiC, XCOPA | Chinese Characters | Sino-Tibetan |
Appendix B Experiment Raw Results
B.1 ICL Modes Evaluation
B.2 Context-Irrelevant Sentence
Appendix C ICL Behavioral Analysis
C.1 Specilized Neuron
Inspired by the universal concept space Wendler et al. (2024); Zhao et al. (2024); Wu et al. (2024), we hypothesize that MLLMs could activate more cross-lingual capabilities by aligning different linguistic representations. To validate the above hypothesis, we seek to find patterns in neuron behavior between ICL modes. Following Tan et al. (2024); Tang et al. (2024); Kojima et al. (2024); Mu et al. (2024); Zhao et al. (2024); Xu et al. (2024), we look at the activations of neurons in the multilayer perceptron (MLP, or Feedforward Network, FFN) modules of the MLLMs.
C.1.1 Identifying Top-Activated Neurons
Each neuron in every MLP layer of the model is assigned a dedicated counter, initialized to . During vanilla evaluation, we monitor the activation of every neuron in the forward pass. Since LLMs typically employ ReLU-like Agarap (2018) activation functions (e.g., SwiGLU Shazeer (2020) for Llama series), a positive activation value can be interpreted as the neuron being “activated”. If a neuron is “activated”, we increment the corresponding counter by , otherwise no action. To ensure balanced input across our three datasets, we curated a COMBINED dataset, see § A.2 for details. After processing the inputs of a single ICL mode, each neuron accumulates an “activated” count. The neurons with the highest counts are identified as specialized neurons attributed to this ICL mode.
C.2 Multilingual-specific Neurons Overlap Most with Native-specific Neurons
| IoU | All Langs | LRL | HRL |
| English– Multilingual– Native | |||
| Native- English | 61.82 | 56.81 | 68.50 |
| Native- Multilingual | 78.84 | 66.19 | 85.10 |
| English- Multilingual | 65.84 | 66.89 | 64.60 |
| Native- English | 70.61 | 66.05 | 81.22 |
| Native- Multilingual | 85.13 | 77.40 | 91.31 |
| English- Multilingual | 78.24 | 79.83 | 77.46 |
| English– Chinese– Native | |||
| Native- English | 61.82 | 56.81 | 68.50 |
| Native- Chinese | 60.58 | 55.35 | 65.73 |
| English- Chinese | 56.52 | 57.94 | 55.72 |
| Native- English | 70.61 | 66.05 | 81.22 |
| Native- Chinese | 73.41 | 69.39 | 82.18 |
| English- Chinese | 77.07 | 78.34 | 76.51 |
We examined the overlaps among the most-activated neurons (specialized neurons) across different ICL modes by calculating the Intersection over Union (IoU) scores. For ICL modes and , with specialized neurons denoted as sets and , their overlap is quantified by Eq. 4:
Prior research Tang et al. (2024); Zhao et al. (2024) has discovered that language-specific neurons are located primarily in the models’ top and bottom layers. Because we want to explain the multilingual reasoning capabilities of a model, we only consider neurons that belong to a certain prefix or suffix of the models’ layers in an effort to restrict our selected neuron sets to be mainly language-specific neurons.
The similarity in performance between Multilingual and Native can be explained by the high overlap in the sets of most-activated neurons. On the other hand, the poorer performance of English can be explained by the low overlap between English most-activated neurons and other ICL modes .
The same experiment was also performed but with Multilingual replaced by HRL Chinese. In this case, the patterns were less obvious and model-specific. This result aligns with our findings in vanilla evaluation that English Non-English HRL Monolingual Multilingual (§ 4.2).
We further split this neuron experiment to either use only LRL or HRL ICL modes when recording neuron activations. We observed that HRL tends to activate similar neurons between Native and Multilingual, whereas LRL tends to activate similar neurons between English and Multilingual.
| MGSM | bn | de | en | es | fr | ja | ru | sw | te | th | zh | LRL Avg | HRL Avg | ALL Avg |
| English | 66.40 | 76.40 | 86.40 | 78.80 | 77.60 | 66.40 | 77.20 | 59.20 | 60.00 | 71.20 | 75.60 | 64.20 | 76.91 | 72.29 |
| French | 67.60 | 77.60 | 86.00 | 77.60 | 75.20 | 70.00 | 79.60 | 56.40 | 59.60 | 68.00 | 70.40 | |||
| Chinese | 66.80 | 77.20 | 82.80 | 79.60 | 74.80 | 68.00 | 76.40 | 55.60 | 59.20 | 70.40 | 72.80 | |||
| Japanese | 66.80 | 78.80 | 83.20 | 76.00 | 74.80 | 70.80 | 76.40 | 56.40 | 56.80 | 68.80 | 68.00 | |||
| Multilingual | 65.20 | 79.60 | 84.40 | 77.60 | 74.00 | 69.60 | 76.80 | 52.40 | 53.60 | 69.60 | 70.00 | |||
| Native | 64.80 | 76.80 | 86.40 | 80.00 | 75.20 | 70.80 | 76.80 | 58.40 | 54.80 | 70.40 | 72.80 | |||
| Transl-LangEn | 63.60 | 73.20 | 86.40 | 78.00 | 74.00 | 60.40 | 77.60 | 69.60 | 60.00 | 46.40 | 58.80 | |||
| Transl-EnLang | 65.20 | 78.80 | 86.40 | 79.60 | 76.00 | 68.40 | 76.40 | 58.80 | 52.80 | 69.20 | 73.60 | |||
| English | 52.40 | 69.20 | 87.20 | 63.60 | 76.80 | 68.00 | 78.80 | 67.20 | 39.60 | 69.20 | 75.20 | 57.10 | 74.11 | 67.93 |
| French | 62.80 | 74.40 | 89.20 | 82.40 | 82.40 | 69.60 | 79.20 | 69.60 | 45.20 | 70.00 | 76.80 | |||
| Chinese | 66.00 | 76.80 | 87.20 | 83.20 | 78.00 | 67.60 | 79.60 | 70.00 | 48.40 | 69.20 | 77.60 | |||
| Japanese | 66.40 | 76.80 | 86.00 | 80.80 | 78.00 | 67.20 | 79.60 | 72.80 | 58.00 | 75.20 | 77.60 | |||
| Multilingual | 68.00 | 76.40 | 88.00 | 84.80 | 80.40 | 69.20 | 80.40 | 68.80 | 55.60 | 71.60 | 75.20 | |||
| Native | 67.20 | 80.40 | 87.20 | 81.20 | 82.40 | 67.20 | 80.00 | 72.40 | 58.00 | 76.40 | 77.60 | |||
| Transl-LangEn | 62.80 | 74.80 | 87.20 | 82.00 | 77.60 | 62.80 | 81.60 | 68.00 | 62.00 | 47.20 | 57.60 | |||
| Transl-EnLang | 67.60 | 80.00 | 87.20 | 82.80 | 83.60 | 70.80 | 80.80 | 70.80 | 59.60 | 76.40 | 79.20 | |||
| English | 57.20 | 74.00 | 90.40 | 82.00 | 80.00 | 67.60 | 80.40 | 20.80 | 23.20 | 73.60 | 79.20 | 43.70 | 79.09 | 66.22 |
| French | 54.80 | 74.00 | 92.00 | 82.00 | 79.60 | 66.80 | 79.20 | 22.40 | 20.80 | 73.60 | 80.40 | |||
| Chinese | 56.00 | 76.40 | 92.00 | 81.20 | 77.60 | 71.20 | 79.60 | 28.80 | 20.40 | 73.60 | 85.20 | |||
| Japanese | 51.20 | 72.80 | 88.80 | 78.80 | 76.00 | 75.20 | 79.20 | 27.20 | 20.80 | 72.40 | 80.80 | |||
| Multilingual | 64.80 | 77.20 | 89.20 | 82.80 | 81.60 | 70.40 | 82.00 | 28.40 | 23.60 | 73.20 | 79.60 | |||
| Native | 72.00 | 83.60 | 90.40 | 82.80 | 79.60 | 75.20 | 83.20 | 32.40 | 40.40 | 76.80 | 85.20 | |||
| Transl-LangEn | 66.00 | 76.00 | 90.40 | 79.60 | 79.20 | 64.00 | 77.60 | 72.40 | 61.20 | 47.60 | 58.40 | |||
| Transl-EnLang | 72.40 | 81.20 | 90.40 | 82.40 | 80.00 | 69.20 | 81.20 | 32.80 | 43.20 | 74.80 | 80.80 | |||
| English | 75.20 | 82.40 | 94.40 | 88.80 | 87.60 | 76.80 | 88.00 | 30.80 | 48.80 | 82.80 | 86.80 | 59.40 | 86.40 | 76.58 |
| French | 74.00 | 85.20 | 92.40 | 90.00 | 85.60 | 78.80 | 86.80 | 33.60 | 51.20 | 82.40 | 84.00 | |||
| Chinese | 77.20 | 85.20 | 94.00 | 88.80 | 86.00 | 81.60 | 88.00 | 31.20 | 50.40 | 81.20 | 86.40 | |||
| Japanese | 74.80 | 84.80 | 94.80 | 90.80 | 88.00 | 81.20 | 88.40 | 32.80 | 50.80 | 83.20 | 85.60 | |||
| Multilingual | 76.40 | 86.80 | 92.80 | 89.60 | 88.00 | 78.40 | 87.20 | 30.40 | 50.00 | 81.20 | 86.00 | |||
| Native | 75.20 | 86.80 | 94.40 | 89.20 | 85.60 | 81.20 | 87.60 | 35.60 | 46.00 | 84.40 | 86.40 | |||
| Transl-LangEn | 70.40 | 79.60 | 94.40 | 83.60 | 82.40 | 64.00 | 81.60 | 72.80 | 62.40 | 48.00 | 64.80 | |||
| Transl-EnLang | 74.80 | 85.60 | 94.40 | 88.80 | 84.40 | 80.40 | 87.20 | 35.20 | 46.80 | 82.80 | 88.00 | |||
| English | 66.80 | 75.60 | 90.80 | 81.20 | 80.00 | 64.40 | 83.60 | 42.00 | 54.00 | 65.20 | 76.80 | 57.00 | 78.91 | 70.95 |
| French | 66.00 | 81.20 | 92.40 | 82.00 | 81.60 | 74.80 | 83.20 | 46.40 | 70.40 | 74.40 | 76.40 | |||
| Chinese | 70.40 | 81.60 | 91.60 | 81.60 | 82.00 | 73.20 | 84.80 | 49.20 | 69.60 | 69.20 | 77.20 | |||
| Japanese | 67.20 | 84.40 | 90.00 | 85.60 | 83.20 | 74.40 | 84.80 | 45.60 | 69.60 | 72.00 | 76.00 | |||
| Multilingual | 67.20 | 82.00 | 90.40 | 82.80 | 77.60 | 71.60 | 83.20 | 50.80 | 66.40 | 68.80 | 80.00 | |||
| Native | 66.40 | 81.20 | 90.80 | 82.00 | 81.60 | 74.40 | 81.60 | 58.00 | 67.60 | 71.20 | 77.20 | |||
| Transl-LangEn | 65.20 | 76.80 | 90.80 | 82.00 | 78.80 | 62.00 | 80.00 | 71.60 | 62.40 | 47.20 | 61.60 | |||
| Transl-EnLang | 68.80 | 83.60 | 90.80 | 80.40 | 82.00 | 74.00 | 85.60 | 57.60 | 66.80 | 73.20 | 77.60 | |||
| English | 39.60 | 75.60 | 82.40 | 81.60 | 72.40 | 67.20 | 77.20 | 20.00 | 17.20 | 40.80 | 71.20 | 29.40 | 75.37 | 58.65 |
| French | 43.60 | 74.40 | 81.60 | 80.00 | 75.20 | 67.60 | 77.60 | 19.60 | 23.20 | 36.40 | 74.00 | |||
| Chinese | 44.00 | 74.40 | 80.40 | 77.60 | 73.20 | 66.00 | 77.20 | 20.00 | 22.40 | 40.40 | 76.40 | |||
| Japanese | 42.80 | 72.40 | 83.20 | 79.20 | 72.80 | 66.80 | 77.20 | 20.40 | 20.00 | 38.40 | 72.00 | |||
| Multilingual | 46.80 | 74.00 | 83.60 | 78.00 | 72.80 | 67.60 | 79.60 | 19.20 | 17.60 | 36.80 | 72.00 | |||
| Native | 44.80 | 77.60 | 82.40 | 79.60 | 75.20 | 66.80 | 79.20 | 19.60 | 22.00 | 44.00 | 76.40 | |||
| Transl-LangEn | 64.40 | 71.20 | 82.40 | 76.40 | 74.40 | 57.60 | 69.20 | 65.20 | 62.80 | 48.00 | 59.60 | |||
| Transl-EnLang | 46.40 | 74.80 | 82.40 | 80.00 | 74.80 | 66.00 | 78.80 | 19.20 | 20.80 | 42.40 | 76.00 | |||
| English | 39.60 | 75.20 | 86.00 | 76.80 | 62.80 | 59.20 | 66.40 | 63.60 | 12.80 | 60.80 | 67.20 | 44.20 | 70.51 | 60.95 |
| French | 34.00 | 74.80 | 86.00 | 84.80 | 79.20 | 66.00 | 67.60 | 63.20 | 15.20 | 55.60 | 74.40 | |||
| Chinese | 30.40 | 78.80 | 83.20 | 78.40 | 77.20 | 69.20 | 77.60 | 67.20 | 15.60 | 62.40 | 73.20 | |||
| Japanese | 27.60 | 78.80 | 85.20 | 82.00 | 72.40 | 73.20 | 74.40 | 69.60 | 14.00 | 61.20 | 76.00 | |||
| Multilingual | 54.40 | 79.60 | 83.60 | 79.20 | 73.60 | 68.00 | 75.20 | 68.00 | 24.40 | 57.20 | 74.40 | |||
| Native | 57.20 | 79.60 | 86.00 | 80.40 | 81.60 | 75.20 | 77.60 | 73.60 | 30.00 | 58.80 | 74.00 | |||
| English | 87.20 | 90.80 | 94.80 | 92.00 | 87.20 | 84.80 | 92.00 | 83.20 | 84.00 | 88.80 | 90.00 | 85.80 | 90.23 | 88.62 |
| French | 87.20 | 90.80 | 94.40 | 92.80 | 89.20 | 82.40 | 90.80 | 85.20 | 84.00 | 88.80 | 87.60 | |||
| Chinese | 86.80 | 90.80 | 93.60 | 93.20 | 90.00 | 83.20 | 91.20 | 85.20 | 84.40 | 90.00 | 90.80 | |||
| Japanese | 85.20 | 89.60 | 94.80 | 92.80 | 86.80 | 86.00 | 92.00 | 82.80 | 81.60 | 88.40 | 87.60 | |||
| Multilingual | 86.40 | 90.00 | 92.80 | 94.00 | 89.20 | 84.00 | 92.40 | 84.00 | 80.40 | 88.00 | 89.20 | |||
| Native | 85.20 | 88.00 | 94.80 | 91.20 | 89.20 | 85.20 | 90.40 | 85.60 | 81.60 | 88.40 | 90.00 | |||
| XCOPA | en | et | ht | id | it | qu | sw | ta | th | tr | vi | zh | LRL AVG | HRL AVG | ALL AVG |
| English | 95.20 | 55.80 | 10.20 | 79.60 | 86.60 | 7.60 | 40.00 | 59.40 | 69.60 | 72.00 | 81.00 | 87.00 | 46.23 | 84.08 | 62.00 |
| Italian | 93.80 | 59.80 | 46.80 | 81.00 | 89.60 | 41.20 | 57.40 | 58.80 | 73.40 | 76.00 | 80.20 | 88.20 | |||
| Chinese | 93.80 | 59.80 | 52.80 | 82.40 | 85.60 | 47.60 | 55.40 | 59.00 | 79.00 | 74.60 | 79.60 | 90.80 | |||
| Multilingual | 94.60 | 57.80 | 51.80 | 81.60 | 87.80 | 46.40 | 60.40 | 60.80 | 79.20 | 78.80 | 80.40 | 88.80 | |||
| Native | 95.20 | 68.60 | 61.60 | 85.00 | 89.60 | 50.40 | 62.00 | 65.80 | 77.80 | 79.80 | 84.80 | 90.80 | |||
| Transl-LangEn | 95.20 | 84.00 | 69.80 | 83.00 | 88.40 | 61.00 | 69.00 | 70.40 | 66.00 | 86.20 | 85.20 | 85.20 | |||
| Transl-EnLang | 95.20 | 69.60 | 62.00 | 87.40 | 89.20 | 52.00 | 62.00 | 63.80 | 78.60 | 79.80 | 84.00 | 89.20 | |||
| English | 95.60 | 65.20 | 27.00 | 84.20 | 88.40 | 26.00 | 52.60 | 62.40 | 73.80 | 76.80 | 84.40 | 89.80 | 55.91 | 86.96 | 68.85 |
| Italian | 95.00 | 62.20 | 28.40 | 86.00 | 92.80 | 32.60 | 61.60 | 70.20 | 76.60 | 78.20 | 85.00 | 90.80 | |||
| Chinese | 95.00 | 65.20 | 48.40 | 83.80 | 88.20 | 32.20 | 58.80 | 68.80 | 76.40 | 78.60 | 86.00 | 91.40 | |||
| Multilingual | 96.00 | 60.40 | 57.20 | 87.60 | 90.80 | 46.80 | 61.60 | 70.80 | 78.60 | 83.00 | 87.40 | 90.80 | |||
| Native | 95.60 | 72.40 | 66.20 | 89.80 | 92.80 | 52.60 | 66.60 | 75.20 | 80.80 | 84.40 | 87.60 | 91.40 | |||
| Transl-LangEn | 95.60 | 87.60 | 74.60 | 88.20 | 89.60 | 68.20 | 73.20 | 74.20 | 72.60 | 88.00 | 86.80 | 87.00 | |||
| Transl-EnLang | 95.60 | 74.40 | 63.40 | 89.00 | 90.00 | 52.20 | 65.20 | 74.00 | 78.40 | 83.60 | 85.80 | 92.00 | |||
| English | 97.00 | 61.80 | 50.60 | 88.00 | 90.20 | 49.80 | 53.20 | 58.40 | 77.80 | 75.40 | 84.40 | 91.00 | 62.29 | 88.32 | 73.13 |
| Italian | 92.20 | 66.40 | 51.60 | 88.40 | 95.40 | 52.20 | 54.40 | 59.80 | 79.80 | 78.20 | 83.80 | 87.20 | |||
| Chinese | 88.60 | 65.00 | 51.80 | 81.20 | 86.00 | 50.20 | 53.00 | 60.00 | 79.00 | 77.60 | 83.60 | 93.60 | |||
| Multilingual | 96.00 | 63.20 | 53.80 | 90.60 | 94.00 | 53.00 | 53.40 | 59.60 | 80.40 | 77.60 | 83.40 | 93.00 | |||
| Native | 97.00 | 71.20 | 54.80 | 93.00 | 95.40 | 51.40 | 60.20 | 63.20 | 83.60 | 81.20 | 89.00 | 93.60 | |||
| Transl-LangEn | 97.00 | 88.40 | 80.00 | 91.00 | 92.00 | 75.60 | 79.80 | 82.80 | 76.20 | 90.20 | 88.40 | 89.40 | |||
| Transl-EnLang | 97.00 | 67.20 | 54.60 | 91.20 | 94.80 | 51.00 | 56.80 | 61.60 | 83.00 | 80.20 | 86.40 | 91.80 | |||
| English | 97.40 | 62.20 | 56.40 | 89.40 | 93.20 | 50.80 | 53.40 | 58.20 | 83.40 | 80.20 | 88.40 | 93.60 | 64.69 | 90.76 | 75.55 |
| Italian | 96.40 | 65.20 | 57.80 | 89.00 | 95.40 | 49.00 | 52.20 | 58.40 | 82.00 | 83.60 | 87.20 | 92.80 | |||
| Chinese | 95.80 | 65.00 | 58.40 | 89.00 | 92.40 | 50.80 | 51.40 | 58.60 | 83.40 | 82.00 | 89.80 | 94.60 | |||
| Multilingual | 97.00 | 65.40 | 56.80 | 91.80 | 94.00 | 49.20 | 49.60 | 59.20 | 83.60 | 84.00 | 88.60 | 92.40 | |||
| Native | 97.40 | 69.60 | 62.80 | 91.40 | 95.40 | 50.60 | 54.00 | 61.40 | 85.40 | 84.40 | 90.00 | 94.60 | |||
| Transl-LangEn | 97.40 | 87.80 | 80.60 | 91.20 | 91.20 | 75.00 | 77.40 | 80.60 | 74.20 | 89.80 | 87.80 | 88.00 | |||
| Transl-EnLang | 97.40 | 67.40 | 61.20 | 91.00 | 96.00 | 46.80 | 53.60 | 60.40 | 85.40 | 82.80 | 88.80 | 95.20 | |||
| English | 96.60 | 57.60 | 58.00 | 83.00 | 93.20 | 50.60 | 51.60 | 73.40 | 64.60 | 73.60 | 80.00 | 91.20 | 62.26 | 87.52 | 72.78 |
| Italian | 96.00 | 57.40 | 55.40 | 82.00 | 95.80 | 48.60 | 53.80 | 73.00 | 64.80 | 73.00 | 81.60 | 90.00 | |||
| Chinese | 96.40 | 58.20 | 52.00 | 83.00 | 92.60 | 49.00 | 53.80 | 70.20 | 66.80 | 71.00 | 81.00 | 91.60 | |||
| Multilingual | 95.80 | 60.20 | 54.00 | 85.80 | 94.40 | 48.80 | 56.00 | 76.20 | 62.80 | 72.60 | 82.60 | 92.20 | |||
| Native | 96.60 | 74.00 | 64.00 | 87.40 | 95.80 | 48.80 | 62.20 | 82.20 | 79.00 | 80.00 | 86.20 | 91.60 | |||
| Transl-LangEn | 96.60 | 88.00 | 75.40 | 87.60 | 89.40 | 70.00 | 74.20 | 77.40 | 74.00 | 88.20 | 85.80 | 84.80 | |||
| Transl-EnLang | 96.60 | 71.40 | 63.00 | 88.00 | 94.40 | 49.20 | 60.60 | 76.40 | 73.20 | 80.60 | 85.40 | 93.00 | |||
| English | 95.40 | 28.20 | 17.60 | 83.80 | 84.20 | 0.00 | 11.20 | 45.20 | 10.00 | 80.40 | 73.60 | 77.60 | 26.54 | 84.28 | 50.60 |
| Italian | 92.00 | 15.40 | 18.40 | 86.40 | 91.40 | 0.20 | 11.20 | 45.80 | 34.60 | 82.20 | 79.60 | 85.60 | |||
| Chinese | 90.60 | 39.40 | 34.80 | 81.00 | 82.80 | 0.00 | 17.60 | 38.60 | 23.00 | 79.00 | 78.40 | 92.20 | |||
| Multilingual | 93.40 | 52.20 | 52.00 | 88.00 | 90.00 | 5.60 | 46.00 | 58.20 | 43.60 | 84.00 | 83.20 | 90.20 | |||
| Native | 95.40 | 54.00 | 56.20 | 88.80 | 91.40 | 53.40 | 53.40 | 69.40 | 62.40 | 85.60 | 86.20 | 92.20 | |||
| Transl-LangEn | 95.40 | 84.40 | 75.20 | 86.60 | 86.80 | 67.40 | 74.40 | 74.60 | 67.80 | 86.40 | 84.20 | 85.60 | |||
| Transl-EnLang | 95.40 | 54.20 | 54.40 | 89.40 | 90.80 | 45.80 | 51.00 | 65.80 | 60.00 | 83.80 | 86.00 | 90.60 | |||
| English | 96.00 | 77.20 | 56.80 | 83.00 | 88.80 | 48.40 | 70.80 | 52.00 | 64.80 | 76.20 | 74.00 | 83.80 | 63.43 | 85.56 | 72.65 |
| Italian | 95.00 | 78.60 | 57.20 | 83.20 | 91.20 | 49.20 | 71.60 | 49.40 | 60.40 | 79.40 | 77.40 | 85.60 | |||
| Chinese | 94.40 | 77.20 | 58.60 | 84.40 | 86.80 | 48.80 | 69.80 | 50.00 | 62.00 | 76.80 | 75.20 | 86.80 | |||
| Multilingual | 93.80 | 75.00 | 56.80 | 87.40 | 89.60 | 49.20 | 71.40 | 48.00 | 62.80 | 85.80 | 75.80 | 86.80 | |||
| Native | 96.00 | 85.20 | 67.60 | 87.20 | 90.80 | 48.00 | 77.20 | 53.20 | 61.40 | 87.60 | 77.60 | 87.60 | |||
| English | 98.60 | 93.20 | 80.00 | 94.20 | 97.60 | 49.80 | 84.20 | 83.40 | 88.20 | 95.20 | 92.80 | 95.60 | 81.66 | 96.24 | 87.73 |
| Italian | 98.80 | 91.60 | 74.40 | 95.00 | 98.20 | 49.20 | 82.60 | 84.40 | 89.00 | 95.60 | 92.80 | 96.40 | |||
| Chinese | 98.00 | 93.40 | 81.00 | 94.20 | 97.80 | 49.80 | 84.00 | 84.40 | 88.80 | 94.80 | 93.40 | 95.00 | |||
| Multilingual | 98.60 | 93.80 | 80.00 | 96.00 | 98.20 | 50.20 | 85.00 | 86.20 | 92.40 | 96.20 | 94.40 | 95.20 | |||
| Native | 98.60 | 94.80 | 89.60 | 95.20 | 98.20 | 52.20 | 87.60 | 88.80 | 93.80 | 95.20 | 95.00 | 95.20 | |||
| XL-WiC | bg | da | de | en | et | fa | fr | hr | it | ja | ko | nl | zh | LRL AVG | HRL AVG | ALL AVG |
| English | 55.13 | 66.15 | 59.49 | 67.44 | 55.38 | 63.33 | 59.49 | 55.64 | 53.85 | 54.62 | 56.67 | 55.90 | 64.10 | 57.37 | 59.74 | 59.01 |
| French | 57.95 | 63.08 | 66.67 | 64.62 | 52.82 | 68.21 | 66.15 | 57.18 | 56.15 | 55.13 | 51.79 | 62.82 | 65.38 | |||
| Chinese | 56.67 | 64.10 | 64.87 | 63.59 | 55.38 | 63.08 | 58.97 | 54.87 | 62.31 | 57.18 | 48.21 | 63.85 | 63.33 | |||
| Japanese | 59.74 | 61.79 | 66.15 | 65.38 | 52.56 | 70.77 | 60.00 | 56.67 | 59.23 | 58.97 | 55.13 | 64.87 | 63.33 | |||
| Multilingual | 57.18 | 60.00 | 62.82 | 64.10 | 53.85 | 69.23 | 60.00 | 56.92 | 58.97 | 58.46 | 57.44 | 62.82 | 58.72 | |||
| Native | 63.59 | 55.90 | 69.23 | 67.44 | 62.05 | 68.72 | 66.15 | 58.21 | 59.49 | 58.97 | 66.67 | 66.41 | 63.33 | |||
| English | 51.54 | 50.77 | 61.28 | 66.92 | 44.62 | 29.74 | 56.67 | 51.79 | 15.90 | 48.46 | 33.85 | 51.28 | 57.44 | 44.42 | 49.17 | 47.71 |
| French | 54.62 | 60.00 | 66.41 | 64.62 | 49.74 | 22.56 | 62.31 | 57.44 | 52.56 | 49.49 | 53.33 | 65.13 | 54.10 | |||
| Chinese | 54.87 | 62.82 | 62.31 | 61.79 | 53.08 | 55.64 | 56.15 | 57.95 | 56.67 | 49.74 | 48.21 | 62.31 | 62.05 | |||
| Japanese | 54.36 | 60.00 | 63.59 | 60.00 | 47.69 | 63.33 | 58.72 | 52.05 | 58.21 | 53.59 | 48.97 | 58.21 | 56.41 | |||
| Multilingual | 57.69 | 60.00 | 65.38 | 62.56 | 55.13 | 58.46 | 57.95 | 56.92 | 55.38 | 54.36 | 52.31 | 64.62 | 58.46 | |||
| Native | 61.79 | 64.10 | 68.46 | 66.92 | 62.05 | 70.51 | 62.31 | 57.18 | 56.15 | 53.59 | 63.08 | 69.49 | 62.05 | |||
| English | 28.21 | 38.46 | 66.67 | 68.46 | 56.92 | 53.85 | 63.33 | 54.87 | 39.23 | 1.28 | 15.90 | 64.36 | 0.77 | 48.46 | 39.83 | 42.49 |
| French | 53.59 | 51.54 | 64.87 | 66.92 | 53.85 | 58.21 | 63.08 | 56.41 | 49.74 | 25.90 | 33.08 | 62.82 | 14.10 | |||
| Chinese | 47.18 | 27.44 | 68.46 | 67.44 | 57.44 | 58.97 | 61.79 | 55.13 | 40.51 | 56.41 | 37.69 | 60.00 | 68.72 | |||
| Japanese | 56.15 | 30.26 | 66.41 | 66.92 | 56.15 | 60.26 | 61.79 | 56.67 | 47.95 | 60.77 | 59.49 | 66.41 | 40.26 | |||
| Multilingual | 53.59 | 60.00 | 65.64 | 65.64 | 58.46 | 58.72 | 60.26 | 54.36 | 56.67 | 58.72 | 66.92 | 67.18 | 57.95 | |||
| Native | 55.13 | 62.05 | 69.49 | 68.46 | 59.23 | 62.05 | 63.08 | 54.62 | 60.51 | 60.77 | 68.21 | 64.87 | 68.72 | |||
| English | 58.97 | 61.28 | 73.33 | 70.51 | 55.64 | 53.59 | 65.90 | 63.08 | 57.95 | 55.64 | 64.10 | 67.69 | 59.74 | 57.82 | 64.02 | 62.11 |
| French | 55.13 | 57.18 | 68.97 | 65.13 | 55.38 | 54.10 | 64.87 | 57.18 | 62.05 | 57.95 | 62.82 | 65.64 | 62.56 | |||
| Chinese | 53.33 | 54.87 | 67.18 | 67.18 | 54.62 | 51.28 | 63.85 | 57.44 | 57.95 | 64.87 | 62.82 | 61.79 | 66.67 | |||
| Japanese | 53.59 | 54.62 | 67.44 | 67.18 | 54.10 | 49.74 | 63.08 | 57.18 | 56.41 | 64.10 | 63.33 | 65.13 | 60.00 | |||
| Multilingual | 57.44 | 60.51 | 69.23 | 68.21 | 54.62 | 55.38 | 61.28 | 56.41 | 57.18 | 64.36 | 66.67 | 67.69 | 64.62 | |||
| Native | 60.00 | 63.33 | 72.56 | 70.51 | 55.90 | 63.33 | 64.87 | 55.64 | 62.31 | 64.10 | 70.77 | 72.82 | 66.67 | |||
| English | 52.56 | 63.85 | 70.51 | 65.64 | 51.79 | 49.23 | 61.03 | 52.56 | 38.97 | 23.08 | 53.59 | 67.44 | 54.10 | 51.54 | 55.36 | 54.18 |
| French | 52.05 | 57.69 | 66.41 | 64.36 | 51.54 | 48.46 | 60.77 | 57.18 | 57.18 | 56.67 | 55.13 | 59.74 | 54.10 | |||
| Chinese | 47.95 | 57.95 | 63.33 | 63.85 | 50.51 | 47.18 | 59.49 | 52.31 | 56.41 | 57.18 | 54.62 | 61.28 | 62.82 | |||
| Japanese | 48.21 | 53.08 | 60.77 | 64.10 | 51.28 | 47.44 | 58.46 | 52.31 | 56.92 | 64.36 | 56.67 | 58.72 | 53.59 | |||
| Multilingual | 51.54 | 56.67 | 67.18 | 61.28 | 52.05 | 49.74 | 61.28 | 51.28 | 54.36 | 57.69 | 56.67 | 58.72 | 59.49 | |||
| Native | 57.44 | 57.69 | 70.00 | 65.64 | 57.69 | 66.92 | 60.77 | 58.72 | 55.13 | 64.36 | 64.87 | 65.90 | 62.82 | |||
| English | 53.08 | 57.44 | 60.51 | 66.41 | 58.46 | 66.41 | 57.44 | 57.18 | 26.41 | 44.62 | 59.74 | 61.79 | 60.51 | 58.78 | 54.99 | 56.15 |
| French | 55.64 | 59.23 | 71.79 | 63.85 | 57.69 | 72.31 | 64.87 | 54.10 | 61.79 | 65.38 | 65.64 | 70.77 | 65.64 | |||
| Chinese | 55.90 | 64.62 | 68.97 | 63.85 | 56.15 | 73.59 | 61.28 | 53.08 | 58.46 | 59.49 | 67.69 | 66.41 | 61.79 | |||
| Japanese | 59.49 | 58.72 | 70.00 | 66.67 | 58.21 | 71.54 | 63.59 | 55.13 | 54.62 | 61.03 | 68.46 | 70.00 | 64.36 | |||
| Multilingual | 58.46 | 58.46 | 69.74 | 63.33 | 56.41 | 69.23 | 66.15 | 55.38 | 61.28 | 65.64 | 68.21 | 71.28 | 63.08 | |||
| Native | 51.54 | 63.08 | 71.54 | 66.41 | 56.92 | 78.97 | 64.87 | 59.49 | 61.03 | 61.03 | 67.95 | 70.51 | 61.79 | |||
| English | 54.36 | 50.77 | 62.31 | 63.59 | 54.62 | 54.10 | 58.46 | 51.79 | 32.82 | 30.77 | 56.67 | 65.13 | 21.03 | 53.72 | 49.06 | 50.49 |
| French | 55.13 | 56.67 | 64.62 | 62.56 | 58.97 | 52.05 | 58.72 | 56.41 | 56.41 | 58.46 | 59.23 | 61.79 | 58.97 | |||
| Chinese | 53.59 | 61.28 | 62.56 | 59.74 | 55.13 | 54.36 | 56.41 | 56.92 | 55.90 | 55.64 | 56.41 | 59.23 | 53.85 | |||
| Japanese | 53.33 | 55.38 | 65.90 | 63.08 | 58.97 | 53.08 | 58.97 | 55.90 | 56.15 | 56.41 | 55.90 | 64.36 | 54.87 | |||
| Multilingual | 52.82 | 60.00 | 66.92 | 59.23 | 60.00 | 54.36 | 60.77 | 56.41 | 53.59 | 56.67 | 61.54 | 63.33 | 59.49 | |||
| Native | 54.62 | 62.56 | 64.87 | 63.59 | 59.49 | 58.46 | 58.21 | 60.26 | 54.36 | 57.95 | 59.74 | 64.87 | 53.85 | |||
| English | 68.72 | 27.95 | 74.36 | 73.33 | 62.56 | 28.46 | 71.79 | 65.64 | 38.21 | 4.10 | 53.59 | 75.64 | 5.38 | 56.35 | 47.15 | 49.98 |
| French | 66.41 | 58.97 | 72.82 | 67.95 | 60.51 | 39.23 | 71.28 | 68.97 | 61.28 | 63.08 | 72.05 | 71.03 | 70.00 | |||
| Chinese | 67.44 | 58.21 | 73.08 | 69.23 | 58.72 | 50.26 | 67.69 | 68.97 | 58.72 | 66.15 | 70.77 | 71.79 | 73.59 | |||
| Japanese | 66.92 | 52.82 | 72.56 | 67.69 | 60.00 | 39.23 | 66.67 | 65.13 | 57.95 | 66.67 | 68.46 | 73.85 | 67.95 | |||
| Multilingual | 66.41 | 65.13 | 72.31 | 68.72 | 62.05 | 64.62 | 68.97 | 67.95 | 65.64 | 64.10 | 71.28 | 73.85 | 67.69 | |||
| Native | 71.54 | 70.26 | 76.15 | 73.33 | 65.13 | 82.82 | 71.79 | 67.95 | 67.95 | 66.15 | 76.41 | 75.64 | 72.05 | |||
| McNemar’s Test | Low-Resource Languages | High-Resource Languages | ||||||||||||
| MGSM ICL Mode | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | ||
| M1 vs M2 Comparison | Level | Wrong | M2 Correct | M2 Wrong | Correct | Level | False | M2 Correct | M2 Wrong | Correct | ||||
| English vs French | 0.85 | 280 | 78 | 91 | 551 | 0.08 | 300 | 104 | 109 | 1237 | ||||
| English vs Chinese | 0.65 | 271 | 87 | 99 | 543 | 1.09 | 295 | 109 | 126 | 1220 | ||||
| English vs Japanese | 2.01 | 278 | 80 | 100 | 542 | 2.58 | 296 | 108 | 134 | 1212 | ||||
| English vs Multilingual | 7.53 | ** | 277 | 81 | 121 | 521 | 1.02 | 302 | 102 | 118 | 1228 | |||
| English vs Native | 2.16 | 276 | 82 | 103 | 539 | 0.00 | 294 | 110 | 109 | 1237 | ||||
| English vs French | 11.75 | *** | 311 | 118 | 70 | 501 | 25.57 | *** | 261 | 192 | 104 | 1193 | ||
| English vs Chinese | 17.39 | *** | 287 | 142 | 79 | 492 | 19.63 | *** | 263 | 190 | 112 | 1185 | ||
| English vs Japanese | 49.92 | *** | 255 | 174 | 64 | 507 | 14.77 | *** | 267 | 186 | 118 | 1179 | ||
| English vs Multilingual | 33.24 | *** | 268 | 161 | 72 | 499 | 24.12 | *** | 248 | 205 | 116 | 1181 | ||
| English vs Native | 50.67 | *** | 246 | 183 | 69 | 502 | 27.57 | *** | 253 | 200 | 107 | 1190 | ||
| English vs French | 0.40 | 505 | 58 | 66 | 371 | 0.00 | 277 | 89 | 88 | 1296 | ||||
| English vs Chinese | 0.53 | 481 | 82 | 72 | 365 | 3.01 | 266 | 100 | 76 | 1308 | ||||
| English vs Japanese | 0.33 | 492 | 71 | 79 | 358 | 0.08 | 264 | 102 | 107 | 1277 | ||||
| English vs Multilingual | 7.36 | ** | 451 | 112 | 74 | 363 | 2.41 | 254 | 112 | 89 | 1295 | |||
| English vs Native | 56.78 | *** | 386 | 177 | 60 | 377 | 20.92 | *** | 232 | 134 | 68 | 1316 | ||
| English vs French | 0.56 | 344 | 62 | 53 | 541 | 0.15 | 186 | 52 | 57 | 1455 | ||||
| English vs Chinese | 0.21 | 343 | 63 | 57 | 537 | 1.07 | 164 | 74 | 61 | 1451 | ||||
| English vs Japanese | 0.69 | 342 | 64 | 54 | 540 | 3.20 | 158 | 80 | 58 | 1454 | ||||
| English vs Multilingual | 0.00 | 342 | 64 | 63 | 531 | 0.71 | 176 | 62 | 52 | 1460 | ||||
| English vs Native | 0.38 | 318 | 88 | 79 | 515 | 1.76 | 166 | 72 | 56 | 1456 | ||||
| English vs French | 22.44 | *** | 278 | 152 | 79 | 491 | 8.98 | ** | 222 | 147 | 99 | 1282 | ||
| English vs Chinese | 23.24 | *** | 271 | 159 | 83 | 487 | 10.82 | ** | 238 | 131 | 82 | 1299 | ||
| English vs Japanese | 17.17 | *** | 274 | 156 | 90 | 480 | 16.32 | *** | 211 | 158 | 93 | 1288 | ||
| English vs Multilingual | 16.93 | *** | 285 | 145 | 82 | 488 | 5.66 | * | 229 | 140 | 102 | 1279 | ||
| English vs Native | 29.80 | *** | 259 | 171 | 83 | 487 | 7.05 | ** | 235 | 134 | 93 | 1288 | ||
| English vs French | 0.81 | 611 | 95 | 82 | 212 | 0.17 | 323 | 108 | 101 | 1218 | ||||
| English vs Chinese | 2.70 | 605 | 101 | 78 | 216 | 0.11 | 317 | 114 | 120 | 1199 | ||||
| English vs Japanese | 0.47 | 614 | 92 | 82 | 212 | 0.31 | 307 | 124 | 134 | 1185 | ||||
| English vs Multilingual | 0.20 | 614 | 92 | 85 | 209 | 0.00 | 296 | 135 | 135 | 1184 | ||||
| English vs Native | 4.62 | * | 586 | 120 | 88 | 206 | 2.47 | 312 | 119 | 95 | 1224 | |||
| English vs French | 1.93 | 455 | 103 | 125 | 317 | 25.29 | *** | 281 | 235 | 137 | 1097 | |||
| English vs Chinese | 0.02 | 438 | 120 | 123 | 319 | 32.82 | *** | 280 | 236 | 126 | 1108 | |||
| English vs Japanese | 0.44 | 449 | 109 | 120 | 322 | 40.56 | *** | 278 | 238 | 117 | 1117 | |||
| English vs Multilingual | 17.81 | *** | 398 | 160 | 92 | 350 | 27.69 | *** | 289 | 227 | 127 | 1107 | ||
| English vs Native | 43.05 | *** | 374 | 184 | 77 | 365 | 65.82 | *** | 264 | 252 | 99 | 1135 | ||
| English vs French | 0.24 | 106 | 36 | 31 | 827 | 0.81 | 136 | 35 | 44 | 1535 | ||||
| English vs Chinese | 0.77 | 106 | 36 | 28 | 830 | 0.05 | 132 | 39 | 36 | 1543 | ||||
| English vs Japanese | 1.69 | 106 | 36 | 49 | 809 | 0.19 | 132 | 39 | 44 | 1535 | ||||
| English vs Multilingual | 1.27 | 108 | 34 | 45 | 813 | 0.01 | 131 | 40 | 40 | 1539 | ||||
| English vs Native | 0.30 | 103 | 39 | 45 | 813 | 1.09 | 139 | 32 | 42 | 1537 | ||||
| McNemar’s Test | Low-Resource Languages | High-Resource Languages | ||||||||||||
| XCOPA ICL Mode | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | ||
| M1 vs M2 Comparison | Level | Wrong | M2 Correct | M2 Wrong | Correct | Level | Wrong | M2 Correct | M2 Wrong | Correct | ||||
| English vs Italian | 274.27 | *** | 1246 | 636 | 166 | 1452 | 8.84 | ** | 287 | 111 | 70 | 2032 | ||
| English vs Chinese | 347.11 | *** | 1177 | 705 | 157 | 1461 | 5.85 | * | 288 | 110 | 76 | 2026 | ||
| English vs Multilingual | 366.08 | *** | 1163 | 719 | 153 | 1465 | 15.76 | *** | 274 | 124 | 68 | 2034 | ||
| English vs Native | 468.19 | *** | 935 | 947 | 210 | 1408 | 45.38 | *** | 240 | 158 | 58 | 2044 | ||
| English vs Italian | 27.32 | *** | 1194 | 349 | 223 | 1734 | 9.27 | ** | 224 | 102 | 62 | 2112 | ||
| English vs Chinese | 74.68 | *** | 1105 | 438 | 216 | 1741 | 0.63 | 241 | 85 | 74 | 2100 | |||
| English vs Multilingual | 157.05 | *** | 961 | 582 | 225 | 1732 | 21.04 | *** | 189 | 137 | 70 | 2104 | ||
| English vs Native | 276.51 | *** | 723 | 820 | 270 | 1687 | 46.05 | *** | 180 | 146 | 50 | 2124 | ||
| English vs Italian | 8.45 | ** | 1084 | 236 | 176 | 2004 | 0.00 | 216 | 76 | 77 | 2131 | |||
| English vs Chinese | 2.59 | 1106 | 214 | 181 | 1999 | 26.05 | *** | 229 | 63 | 136 | 2072 | |||
| English vs Multilingual | 6.11 | * | 1063 | 257 | 203 | 1977 | 14.73 | *** | 193 | 99 | 51 | 2157 | ||
| English vs Native | 40.18 | *** | 796 | 524 | 337 | 1843 | 47.82 | *** | 157 | 135 | 42 | 2166 | ||
| English vs Italian | 0.03 | 937 | 299 | 304 | 1960 | 1.82 | 152 | 79 | 62 | 2207 | ||||
| English vs Chinese | 1.03 | 990 | 246 | 223 | 2041 | 0.01 | 167 | 64 | 64 | 2205 | ||||
| English vs Multilingual | 0.00 | 931 | 305 | 307 | 1957 | 4.60 | * | 144 | 87 | 60 | 2209 | |||
| English vs Native | 11.49 | *** | 713 | 523 | 418 | 1846 | 13.65 | *** | 130 | 101 | 54 | 2215 | ||
| English vs Italian | 0.05 | 1054 | 267 | 273 | 1906 | 0.06 | 234 | 78 | 82 | 2106 | ||||
| English vs Chinese | 0.93 | 1048 | 273 | 297 | 1882 | 1.06 | 227 | 85 | 100 | 2088 | ||||
| English vs Multilingual | 0.72 | 943 | 378 | 354 | 1825 | 1.22 | 212 | 100 | 84 | 2104 | ||||
| English vs Native | 88.46 | *** | 654 | 667 | 364 | 1815 | 23.24 | *** | 178 | 134 | 65 | 2123 | ||
| English vs Italian | 18.11 | *** | 2268 | 303 | 206 | 723 | 28.96 | *** | 242 | 151 | 70 | 2037 | ||
| English vs Chinese | 100.08 | *** | 2194 | 377 | 147 | 782 | 1.51 | 250 | 143 | 122 | 1985 | |||
| English vs Multilingual | 637.99 | *** | 1714 | 857 | 82 | 847 | 63.44 | *** | 219 | 174 | 53 | 2054 | ||
| English vs Native | 1003.90 | *** | 1176 | 1395 | 149 | 780 | 104.47 | *** | 192 | 201 | 41 | 2066 | ||
| English vs Italian | 0.00 | 890 | 390 | 391 | 1829 | 3.89 | * | 213 | 148 | 115 | 2024 | |||
| English vs Chinese | 0.14 | 860 | 420 | 432 | 1788 | 0.11 | 199 | 162 | 155 | 1984 | ||||
| English vs Multilingual | 0.67 | 864 | 416 | 441 | 1779 | 22.46 | *** | 190 | 171 | 93 | 2046 | |||
| English vs Native | 15.10 | *** | 655 | 625 | 494 | 1726 | 39.95 | *** | 170 | 191 | 85 | 2054 | ||
| English vs Italian | 5.19 | * | 529 | 113 | 151 | 2707 | 2.82 | 57 | 37 | 23 | 2383 | |||
| English vs Chinese | 1.00 | 521 | 121 | 105 | 2753 | 0.61 | 68 | 26 | 33 | 2373 | ||||
| English vs Multilingual | 9.63 | ** | 481 | 161 | 109 | 2749 | 3.32 | 57 | 37 | 22 | 2384 | |||
| English vs Native | 51.49 | *** | 348 | 294 | 143 | 2715 | 0.40 | 60 | 34 | 28 | 2378 | |||
| McNemar’s Test | Low-Resource Languages | High-Resource Languages | ||||||||||||
| XL-WiC ICL Mode | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | ||
| M1 vs M2 Comparison | Level | Wrong | M2 Correct | M2 Wrong | Correct | Level | Wrong | M2 Correct | M2 Wrong | Correct | ||||
| English vs French | 1.85 | 483 | 182 | 156 | 739 | 3.70 | 991 | 422 | 367 | 1730 | ||||
| English vs Chinese | 0.00 | 469 | 196 | 194 | 701 | 1.38 | 1000 | 413 | 379 | 1718 | ||||
| English vs Japanese | 3.67 | 438 | 227 | 187 | 708 | 4.59 | * | 905 | 508 | 441 | 1656 | |||
| English vs Multilingual | 1.89 | 427 | 238 | 208 | 687 | 0.45 | 917 | 496 | 474 | 1623 | ||||
| English vs Native | 13.85 | *** | 334 | 331 | 241 | 654 | 19.84 | *** | 856 | 557 | 417 | 1680 | ||
| English vs French | 1.47 | 641 | 226 | 200 | 493 | 113.05 | *** | 1130 | 654 | 321 | 1405 | |||
| English vs Chinese | 66.13 | *** | 563 | 304 | 133 | 560 | 94.72 | *** | 1125 | 659 | 349 | 1377 | ||
| English vs Japanese | 50.57 | *** | 555 | 312 | 157 | 536 | 96.34 | *** | 1195 | 589 | 296 | 1430 | ||
| English vs Multilingual | 82.26 | *** | 535 | 332 | 135 | 558 | 94.44 | *** | 985 | 799 | 454 | 1272 | ||
| English vs Native | 134.15 | *** | 416 | 451 | 163 | 530 | 197.07 | *** | 956 | 828 | 346 | 1380 | ||
| English vs French | 26.64 | *** | 526 | 278 | 168 | 588 | 99.87 | *** | 1559 | 553 | 266 | 1132 | ||
| English vs Chinese | 25.96 | *** | 578 | 226 | 129 | 627 | 274.42 | *** | 1392 | 720 | 213 | 1185 | ||
| English vs Japanese | 34.50 | *** | 463 | 341 | 203 | 553 | 311.88 | *** | 1347 | 765 | 212 | 1186 | ||
| English vs Multilingual | 20.45 | *** | 385 | 419 | 297 | 459 | 468.48 | *** | 1070 | 1042 | 260 | 1138 | ||
| English vs Native | 35.09 | *** | 436 | 368 | 223 | 533 | 645.95 | *** | 1059 | 1053 | 165 | 1233 | ||
| English vs French | 5.71 | * | 563 | 95 | 132 | 770 | 2.29 | 1028 | 235 | 270 | 1977 | |||
| English vs Chinese | 13.46 | *** | 570 | 88 | 145 | 757 | 1.80 | 959 | 304 | 339 | 1908 | |||
| English vs Japanese | 17.89 | *** | 576 | 82 | 147 | 755 | 5.40 | * | 991 | 272 | 330 | 1917 | ||
| English vs Multilingual | 3.10 | 546 | 112 | 141 | 761 | 0.30 | 973 | 290 | 276 | 1971 | ||||
| English vs Native | 0.45 | 462 | 196 | 182 | 720 | 25.30 | *** | 902 | 361 | 237 | 2010 | |||
| English vs French | 0.95 | 686 | 70 | 58 | 746 | 25.69 | *** | 1167 | 400 | 268 | 1675 | |||
| English vs Chinese | 7.17 | ** | 705 | 51 | 83 | 721 | 28.59 | *** | 1098 | 469 | 318 | 1625 | ||
| English vs Japanese | 5.08 | * | 703 | 53 | 80 | 724 | 16.60 | *** | 1147 | 420 | 309 | 1634 | ||
| English vs Multilingual | 0.17 | 684 | 72 | 78 | 726 | 28.68 | *** | 1176 | 391 | 254 | 1689 | |||
| English vs Native | 44.12 | *** | 485 | 271 | 136 | 668 | 87.70 | *** | 1023 | 544 | 275 | 1668 | ||
| English vs French | 0.50 | 346 | 297 | 279 | 638 | 118.02 | *** | 829 | 751 | 384 | 1546 | |||
| English vs Chinese | 0.31 | 361 | 282 | 268 | 649 | 89.50 | *** | 919 | 661 | 358 | 1572 | |||
| English vs Japanese | 2.26 | 354 | 289 | 253 | 664 | 91.35 | *** | 855 | 725 | 403 | 1527 | |||
| English vs Multilingual | 0.43 | 337 | 306 | 289 | 628 | 109.78 | *** | 813 | 767 | 407 | 1523 | |||
| English vs Native | 3.20 | 304 | 339 | 293 | 624 | 133.37 | *** | 904 | 676 | 312 | 1618 | |||
| MGSM CIS | bn | de | en | es | fr | ja | ru | sw | te | th | zh | LRL AVG | HRL AVG | ALL AVG |
| English CIS-En | 64.40 | 75.20 | 84.80 | 78.40 | 74.00 | 69.20 | 74.00 | 56.80 | 49.20 | 73.60 | 72.40 | 61.00 | 75.43 | 70.18 |
| English CIS-Fr | 66.00 | 77.20 | 85.20 | 80.40 | 76.40 | 67.20 | 75.60 | 54.00 | 58.40 | 68.00 | 70.80 | 61.60 | 76.11 | 70.84 |
| English CIS-Ja | 66.40 | 75.60 | 83.20 | 78.80 | 74.80 | 66.80 | 77.20 | 53.20 | 53.60 | 72.00 | 70.80 | 61.30 | 75.31 | 70.22 |
| English CIS-Zh | 63.60 | 76.80 | 83.20 | 77.20 | 76.40 | 69.20 | 77.60 | 52.40 | 52.40 | 68.80 | 71.60 | 59.30 | 76.00 | 69.93 |
| English CIS-Multi | 65.60 | 75.20 | 83.20 | 78.40 | 77.20 | 66.40 | 76.40 | 51.60 | 55.60 | 69.20 | 70.40 | 60.50 | 75.31 | 69.93 |
| English CIS-En | 52.40 | 70.80 | 88.40 | 70.80 | 75.20 | 70.40 | 74.00 | 67.60 | 38.40 | 65.20 | 73.20 | 55.90 | 74.69 | 67.85 |
| English CIS-Fr | 52.80 | 74.40 | 89.20 | 79.60 | 76.80 | 69.20 | 78.80 | 60.00 | 37.60 | 58.00 | 76.40 | 52.10 | 77.77 | 68.44 |
| English CIS-Ja | 61.20 | 76.00 | 87.60 | 78.40 | 75.20 | 70.80 | 78.00 | 66.40 | 41.20 | 66.40 | 75.60 | 58.80 | 77.37 | 70.62 |
| English CIS-Zh | 58.00 | 72.40 | 89.60 | 80.80 | 76.00 | 66.00 | 77.20 | 69.60 | 38.80 | 53.60 | 76.00 | 55.00 | 76.86 | 68.91 |
| English CIS-Multi | 62.00 | 74.80 | 88.80 | 81.60 | 78.00 | 69.20 | 79.60 | 68.80 | 46.40 | 72.80 | 78.80 | 62.50 | 78.69 | 72.80 |
| English CIS-En | 54.40 | 74.00 | 89.60 | 82.00 | 78.80 | 69.60 | 79.20 | 24.80 | 19.20 | 73.60 | 81.20 | 43.00 | 79.20 | 66.04 |
| English CIS-Fr | 55.20 | 75.20 | 90.00 | 84.00 | 72.80 | 68.00 | 80.80 | 26.40 | 18.00 | 74.40 | 82.00 | 43.50 | 78.97 | 66.07 |
| English CIS-Ja | 56.00 | 75.20 | 90.40 | 78.00 | 76.80 | 69.20 | 79.20 | 24.80 | 20.40 | 74.00 | 80.00 | 43.80 | 78.40 | 65.82 |
| English CIS-Zh | 55.60 | 74.80 | 90.80 | 79.20 | 76.80 | 67.60 | 80.80 | 21.60 | 18.40 | 75.20 | 80.40 | 42.70 | 78.63 | 65.56 |
| English CIS-Multi | 54.80 | 76.40 | 89.60 | 82.00 | 76.80 | 69.20 | 80.00 | 24.40 | 18.00 | 73.20 | 81.20 | 42.60 | 79.31 | 65.96 |
| English CIS-En | 75.20 | 86.40 | 92.40 | 88.00 | 87.60 | 80.80 | 88.00 | 31.60 | 48.00 | 82.80 | 85.20 | 59.40 | 86.91 | 76.91 |
| English CIS-Fr | 75.60 | 85.20 | 94.00 | 89.60 | 86.00 | 78.80 | 87.20 | 31.60 | 46.40 | 84.00 | 82.80 | 59.40 | 86.23 | 76.47 |
| English CIS-Ja | 76.80 | 86.40 | 92.40 | 90.40 | 85.60 | 81.20 | 86.00 | 29.60 | 50.40 | 83.60 | 82.00 | 60.10 | 86.29 | 76.76 |
| English CIS-Zh | 78.00 | 88.40 | 93.60 | 89.20 | 88.00 | 80.40 | 90.00 | 29.20 | 52.40 | 84.00 | 82.80 | 60.90 | 87.49 | 77.82 |
| English CIS-Multi | 76.00 | 86.40 | 94.00 | 90.40 | 84.80 | 80.80 | 86.80 | 30.00 | 51.60 | 80.40 | 82.80 | 59.50 | 86.57 | 76.73 |
| English CIS-En | 61.20 | 80.00 | 92.00 | 82.00 | 82.40 | 71.60 | 82.00 | 48.40 | 62.40 | 70.80 | 78.00 | 60.70 | 81.14 | 73.71 |
| English CIS-Fr | 64.80 | 83.20 | 90.40 | 84.00 | 82.80 | 72.00 | 84.40 | 46.40 | 63.20 | 71.20 | 78.00 | 61.40 | 82.11 | 74.58 |
| English CIS-Ja | 67.20 | 82.80 | 92.40 | 81.20 | 83.60 | 73.60 | 84.80 | 43.20 | 63.20 | 67.20 | 76.80 | 60.20 | 82.17 | 74.18 |
| English CIS-Zh | 62.80 | 82.00 | 91.60 | 80.40 | 83.20 | 70.40 | 86.00 | 44.40 | 65.60 | 69.20 | 80.40 | 60.50 | 82.00 | 74.18 |
| English CIS-Multi | 72.00 | 81.60 | 92.00 | 84.00 | 83.60 | 73.60 | 86.40 | 47.20 | 67.20 | 73.20 | 79.20 | 64.90 | 82.91 | 76.36 |
| English CIS-En | 36.00 | 76.40 | 82.40 | 83.20 | 74.80 | 69.20 | 77.60 | 24.00 | 15.60 | 35.60 | 70.80 | 27.80 | 76.34 | 58.69 |
| English CIS-Fr | 42.00 | 72.80 | 83.60 | 79.20 | 73.20 | 72.00 | 76.80 | 20.80 | 17.20 | 34.80 | 72.40 | 28.70 | 75.71 | 58.62 |
| English CIS-Ja | 40.80 | 76.40 | 82.00 | 79.60 | 74.80 | 68.80 | 76.00 | 23.20 | 18.00 | 35.20 | 73.20 | 29.30 | 75.83 | 58.91 |
| English CIS-Zh | 38.40 | 74.40 | 82.00 | 79.60 | 75.20 | 72.00 | 80.00 | 20.40 | 18.40 | 37.20 | 72.40 | 28.60 | 76.51 | 59.09 |
| English CIS-Multi | 39.60 | 76.40 | 81.60 | 81.60 | 73.60 | 71.60 | 77.60 | 22.80 | 18.80 | 30.40 | 73.20 | 27.90 | 76.51 | 58.84 |
| XCOPA CIS | en | et | ht | id | it | qu | sw | ta | th | tr | vi | zh | LRL AVG | HRL AVG | ALL AVG |
| English CIS-En | 95.40 | 56.60 | 16.00 | 80.20 | 84.80 | 18.60 | 39.00 | 58.40 | 69.80 | 70.80 | 80.20 | 87.40 | 48.37 | 83.72 | 63.10 |
| English CIS-Fr | 95.40 | 57.00 | 47.00 | 80.40 | 85.80 | 42.20 | 52.40 | 57.80 | 72.40 | 74.20 | 80.60 | 86.60 | 58.49 | 84.48 | 69.32 |
| English CIS-Ja | 95.00 | 57.20 | 39.60 | 82.20 | 86.60 | 46.40 | 56.00 | 57.00 | 73.20 | 74.00 | 81.40 | 85.60 | 58.69 | 84.68 | 69.52 |
| English CIS-Zh | 95.00 | 58.40 | 35.80 | 82.00 | 86.00 | 43.00 | 56.80 | 57.00 | 73.80 | 73.60 | 81.40 | 87.60 | 58.03 | 84.84 | 69.20 |
| English CIS-Multi | 95.60 | 58.40 | 51.00 | 81.60 | 86.80 | 49.00 | 59.20 | 56.40 | 73.60 | 74.60 | 80.40 | 88.00 | 61.14 | 85.32 | 71.22 |
| English CIS-En | 96.00 | 64.00 | 25.00 | 84.80 | 88.60 | 27.60 | 51.60 | 62.20 | 72.80 | 76.80 | 85.00 | 90.00 | 55.46 | 87.24 | 68.70 |
| English CIS-Fr | 95.20 | 65.80 | 23.00 | 85.00 | 88.20 | 42.80 | 53.20 | 69.80 | 76.00 | 78.00 | 85.20 | 89.20 | 59.40 | 87.12 | 70.95 |
| English CIS-Ja | 96.40 | 65.80 | 53.60 | 86.60 | 89.00 | 46.60 | 61.40 | 68.60 | 77.00 | 79.20 | 86.60 | 88.40 | 65.66 | 87.92 | 74.93 |
| English CIS-Zh | 95.80 | 65.40 | 52.00 | 86.80 | 88.80 | 47.40 | 58.40 | 68.40 | 76.00 | 78.40 | 85.00 | 89.00 | 64.66 | 87.76 | 74.28 |
| English CIS-Multi | 96.40 | 63.20 | 51.20 | 86.80 | 89.80 | 48.80 | 59.80 | 69.20 | 75.60 | 80.80 | 85.40 | 89.40 | 64.74 | 88.64 | 74.70 |
| English CIS-En | 97.00 | 62.00 | 51.40 | 87.80 | 90.20 | 51.00 | 52.40 | 56.40 | 79.00 | 76.20 | 84.00 | 90.20 | 62.31 | 88.28 | 73.13 |
| English CIS-Fr | 97.00 | 62.60 | 49.80 | 87.20 | 90.20 | 50.20 | 54.00 | 57.20 | 79.20 | 74.40 | 83.20 | 90.60 | 62.31 | 87.88 | 72.97 |
| English CIS-Ja | 97.00 | 62.60 | 51.40 | 88.40 | 89.60 | 50.80 | 54.40 | 58.40 | 80.20 | 75.00 | 83.80 | 91.80 | 63.09 | 88.36 | 73.62 |
| English CIS-Zh | 97.20 | 60.80 | 51.00 | 86.80 | 88.80 | 50.60 | 53.60 | 57.80 | 80.20 | 76.20 | 83.40 | 92.80 | 62.49 | 88.36 | 73.27 |
| English CIS-Multi | 97.20 | 62.20 | 52.20 | 87.80 | 90.80 | 51.80 | 52.20 | 56.80 | 78.80 | 74.60 | 83.60 | 90.40 | 62.51 | 88.16 | 73.20 |
| English CIS-En | 97.00 | 63.00 | 56.60 | 89.40 | 91.00 | 50.20 | 52.40 | 56.00 | 83.40 | 78.20 | 87.80 | 94.20 | 64.20 | 89.96 | 74.93 |
| English CIS-Fr | 96.40 | 63.60 | 58.80 | 90.00 | 93.00 | 49.00 | 51.00 | 56.80 | 82.20 | 79.20 | 87.40 | 94.00 | 64.11 | 90.52 | 75.12 |
| English CIS-Ja | 96.60 | 63.60 | 60.00 | 90.00 | 92.60 | 50.00 | 48.80 | 56.00 | 82.80 | 78.80 | 88.20 | 93.80 | 64.20 | 90.36 | 75.10 |
| English CIS-Zh | 97.20 | 64.00 | 59.80 | 89.80 | 92.60 | 50.20 | 52.00 | 58.00 | 84.00 | 79.20 | 88.40 | 93.60 | 65.20 | 90.48 | 75.73 |
| English CIS-Multi | 96.60 | 64.60 | 61.20 | 90.20 | 93.00 | 52.00 | 48.40 | 59.20 | 82.20 | 80.00 | 87.80 | 93.60 | 65.06 | 90.68 | 75.73 |
| English CIS-En | 96.40 | 54.40 | 55.20 | 79.80 | 90.80 | 49.60 | 54.20 | 69.80 | 63.80 | 70.40 | 80.00 | 90.00 | 61.00 | 85.48 | 71.20 |
| English CIS-Fr | 96.00 | 55.80 | 57.40 | 80.20 | 90.00 | 49.40 | 53.80 | 71.20 | 61.20 | 69.60 | 79.80 | 88.60 | 61.23 | 84.88 | 71.08 |
| English CIS-Ja | 95.60 | 58.40 | 54.60 | 82.40 | 91.20 | 51.20 | 54.40 | 72.00 | 61.40 | 69.20 | 80.60 | 88.20 | 61.80 | 85.32 | 71.60 |
| English CIS-Zh | 95.60 | 56.40 | 53.20 | 82.40 | 91.40 | 50.40 | 54.80 | 72.20 | 61.00 | 71.00 | 80.00 | 90.00 | 61.14 | 86.08 | 71.53 |
| English CIS-Multi | 96.20 | 57.20 | 56.20 | 83.40 | 92.20 | 50.60 | 55.40 | 73.40 | 61.80 | 72.80 | 81.00 | 90.20 | 62.23 | 86.96 | 72.53 |
| English CIS-En | 93.80 | 16.40 | 8.60 | 82.80 | 85.20 | 1.80 | 3.20 | 40.00 | 17.60 | 81.40 | 74.80 | 74.80 | 23.20 | 83.60 | 48.37 |
| English CIS-Fr | 94.20 | 27.80 | 13.00 | 84.00 | 87.80 | 0.60 | 5.00 | 47.00 | 28.20 | 81.60 | 79.40 | 79.40 | 28.71 | 85.40 | 52.33 |
| English CIS-Ja | 94.00 | 31.60 | 12.80 | 85.00 | 87.40 | 0.80 | 5.40 | 62.00 | 40.60 | 81.40 | 78.20 | 83.60 | 33.06 | 86.28 | 55.23 |
| English CIS-Zh | 94.80 | 28.60 | 12.00 | 84.80 | 87.20 | 0.40 | 5.40 | 43.60 | 21.80 | 81.60 | 76.80 | 83.60 | 26.94 | 86.40 | 51.72 |
| English CIS-Multi | 94.60 | 36.40 | 18.80 | 86.40 | 89.00 | 0.80 | 12.80 | 49.80 | 37.00 | 82.80 | 79.60 | 83.60 | 33.60 | 87.28 | 55.97 |
| XL-WiC CIS | bg | da | de | en | et | fa | fr | hr | it | ja | ko | nl | zh | LRL AVG | HRL AVG | ALL AVG |
| English CIS-En | 55.13 | 65.90 | 61.54 | 65.64 | 53.85 | 64.36 | 59.23 | 53.59 | 52.05 | 52.82 | 54.62 | 56.92 | 63.59 | 56.73 | 59.15 | 58.40 |
| English CIS-Fr | 56.15 | 67.18 | 63.08 | 64.62 | 54.62 | 65.13 | 60.00 | 54.10 | 53.85 | 52.82 | 55.64 | 60.00 | 60.26 | 57.50 | 59.72 | 59.03 |
| English CIS-Ja | 56.67 | 67.95 | 62.05 | 64.87 | 54.10 | 66.67 | 60.77 | 55.13 | 56.67 | 53.85 | 53.85 | 59.23 | 62.05 | 58.14 | 60.14 | 59.53 |
| English CIS-Zh | 57.44 | 68.72 | 63.33 | 66.15 | 54.36 | 65.64 | 58.97 | 55.64 | 54.10 | 53.33 | 54.87 | 60.26 | 58.97 | 58.27 | 59.86 | 59.37 |
| English CIS-Multi | 56.67 | 65.90 | 62.31 | 64.36 | 54.62 | 65.90 | 60.77 | 54.10 | 54.62 | 51.79 | 52.31 | 59.49 | 61.03 | 57.82 | 59.17 | 58.76 |
| English CIS-En | 53.08 | 52.05 | 62.05 | 64.62 | 41.79 | 44.36 | 54.36 | 52.31 | 31.03 | 48.46 | 52.56 | 54.36 | 57.95 | 47.88 | 53.05 | 51.46 |
| English CIS-Fr | 55.13 | 58.97 | 63.08 | 64.87 | 53.08 | 47.95 | 60.77 | 55.13 | 53.59 | 50.77 | 55.38 | 60.51 | 58.97 | 52.82 | 58.55 | 56.79 |
| English CIS-Ja | 55.13 | 60.77 | 64.87 | 65.13 | 53.33 | 61.79 | 59.74 | 53.59 | 54.62 | 50.77 | 53.85 | 62.05 | 54.87 | 55.96 | 58.52 | 57.73 |
| English CIS-Zh | 54.62 | 58.97 | 64.36 | 65.64 | 53.59 | 58.46 | 61.28 | 52.05 | 55.13 | 51.28 | 54.36 | 61.28 | 57.95 | 54.68 | 58.92 | 57.61 |
| English CIS-Multi | 55.38 | 62.56 | 62.56 | 64.62 | 52.31 | 61.03 | 61.28 | 55.38 | 52.82 | 51.79 | 51.03 | 64.10 | 58.46 | 56.03 | 58.80 | 57.95 |
| English CIS-En | 39.74 | 40.26 | 61.54 | 66.15 | 55.38 | 54.36 | 61.28 | 54.10 | 37.95 | 8.21 | 13.85 | 55.90 | 0.26 | 50.90 | 38.38 | 42.23 |
| English CIS-Fr | 53.33 | 43.59 | 63.85 | 65.90 | 52.05 | 55.64 | 62.05 | 53.59 | 40.26 | 18.21 | 15.90 | 56.41 | 3.33 | 53.65 | 41.05 | 44.93 |
| English CIS-Ja | 56.41 | 50.77 | 62.05 | 65.64 | 53.33 | 61.28 | 60.26 | 53.85 | 43.85 | 48.46 | 34.10 | 60.26 | 15.13 | 56.22 | 48.95 | 51.18 |
| English CIS-Zh | 56.41 | 50.00 | 62.56 | 65.38 | 54.87 | 61.79 | 61.03 | 54.10 | 43.59 | 37.69 | 30.77 | 59.49 | 14.10 | 56.79 | 47.18 | 50.14 |
| English CIS-Multi | 55.13 | 46.67 | 62.82 | 64.10 | 52.05 | 59.49 | 61.03 | 53.08 | 40.26 | 26.67 | 20.00 | 55.64 | 6.92 | 54.94 | 42.68 | 46.45 |
| English CIS-En | 59.49 | 61.79 | 72.05 | 73.33 | 57.44 | 60.51 | 66.67 | 58.72 | 60.77 | 52.56 | 64.87 | 74.62 | 63.59 | 59.04 | 65.58 | 63.57 |
| English CIS-Fr | 55.90 | 62.56 | 71.03 | 72.05 | 57.18 | 62.31 | 64.36 | 60.77 | 59.49 | 57.95 | 66.67 | 71.28 | 64.10 | 59.04 | 65.50 | 63.51 |
| English CIS-Ja | 57.95 | 62.31 | 72.05 | 71.54 | 58.46 | 61.28 | 64.10 | 59.74 | 58.72 | 65.13 | 64.62 | 73.59 | 65.64 | 59.36 | 66.41 | 64.24 |
| English CIS-Zh | 56.92 | 61.54 | 72.56 | 71.28 | 57.44 | 61.54 | 64.36 | 61.03 | 58.72 | 65.13 | 65.13 | 71.79 | 65.64 | 59.23 | 66.24 | 64.08 |
| English CIS-Multi | 57.18 | 62.31 | 71.03 | 72.82 | 57.95 | 63.59 | 65.13 | 58.72 | 57.95 | 64.10 | 64.87 | 73.08 | 65.90 | 59.36 | 66.35 | 64.20 |
| English CIS-En | 48.97 | 62.56 | 69.74 | 66.67 | 51.79 | 47.95 | 58.21 | 49.74 | 34.10 | 38.46 | 52.56 | 67.44 | 53.85 | 49.62 | 55.95 | 54.00 |
| English CIS-Fr | 48.72 | 62.31 | 69.74 | 65.90 | 51.79 | 48.97 | 59.23 | 52.82 | 42.56 | 52.82 | 51.79 | 62.82 | 55.13 | 50.58 | 58.03 | 55.74 |
| English CIS-Ja | 48.21 | 62.05 | 71.28 | 66.15 | 51.03 | 47.95 | 58.72 | 52.31 | 37.18 | 57.69 | 53.85 | 65.13 | 54.62 | 49.87 | 58.52 | 55.86 |
| English CIS-Zh | 48.46 | 60.00 | 69.49 | 65.64 | 51.54 | 49.74 | 58.46 | 50.51 | 37.18 | 54.36 | 53.85 | 65.13 | 57.18 | 50.06 | 57.92 | 55.50 |
| English CIS-Multi | 48.46 | 58.97 | 69.74 | 66.15 | 52.31 | 47.95 | 61.28 | 52.05 | 39.23 | 54.10 | 54.36 | 65.90 | 56.15 | 50.19 | 58.43 | 55.90 |
| English CIS-En | 52.31 | 55.64 | 57.44 | 65.13 | 56.92 | 64.62 | 54.10 | 57.44 | 19.49 | 49.74 | 54.62 | 55.38 | 57.95 | 57.82 | 52.17 | 53.91 |
| English CIS-Fr | 56.92 | 60.00 | 64.10 | 65.13 | 59.49 | 73.59 | 59.74 | 56.15 | 38.97 | 63.33 | 64.62 | 63.59 | 59.74 | 61.54 | 59.91 | 60.41 |
| English CIS-Ja | 55.90 | 58.21 | 62.82 | 65.38 | 61.28 | 71.54 | 54.62 | 55.38 | 28.21 | 64.36 | 64.10 | 61.03 | 58.46 | 61.03 | 57.46 | 58.56 |
| English CIS-Zh | 56.15 | 57.95 | 63.08 | 65.13 | 60.51 | 70.77 | 54.36 | 54.87 | 29.23 | 63.08 | 64.62 | 61.03 | 60.51 | 60.58 | 57.66 | 58.56 |
| English CIS-Multi | 57.95 | 60.77 | 64.36 | 65.64 | 59.49 | 74.36 | 56.41 | 57.18 | 39.49 | 62.31 | 65.64 | 64.87 | 58.21 | 62.24 | 59.74 | 60.51 |
| McNemar’s Test | Low-Resource Languages | High-Resource Languages | ||||||||||||
| MGSM CIS Mode | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | ||
| M1 vs M2 Comparison | Level | Wrong | M2 Correct | M2 Wrong | Correct | Level | False | M2 Correct | M2 Wrong | Correct | ||||
| English CIS-En vs English CIS-Zh | 1.57 | 2.10E-01 | 317 | 73 | 90 | 520 | 0.47 | 4.93E-01 | 339 | 91 | 81 | 1239 | ||
| English CIS-En vs English CIS-Fr | 0.16 | 6.85E-01 | 311 | 79 | 73 | 537 | 0.60 | 4.39E-01 | 323 | 107 | 95 | 1225 | ||
| English CIS-En vs English CIS-Ja | 0.03 | 8.72E-01 | 311 | 79 | 76 | 534 | 0.01 | 9.40E-01 | 343 | 87 | 89 | 1231 | ||
| English CIS-En vs English CIS-Multi | 0.08 | 7.72E-01 | 297 | 93 | 98 | 512 | 0.01 | 9.41E-01 | 341 | 89 | 91 | 1229 | ||
| Multilingual CIS-Multi vs English CIS-Multi | 0.06 | 8.00E-01 | 327 | 72 | 68 | 533 | 0.16 | 6.93E-01 | 355 | 83 | 77 | 1235 | ||
| English CIS-En vs English CIS-Zh | 0.27 | 6.02E-01 | 328 | 113 | 122 | 437 | 5.35 | 2.08E-02 | * | 296 | 147 | 109 | 1198 | |
| English CIS-En vs English CIS-Fr | 5.90 | 1.51E-02 | * | 344 | 97 | 135 | 424 | 10.48 | 1.21E-03 | ** | 282 | 161 | 107 | 1200 |
| English CIS-En vs English CIS-Ja | 3.79 | 5.16E-02 | 323 | 118 | 89 | 470 | 7.69 | 5.54E-03 | ** | 282 | 161 | 114 | 1193 | |
| English CIS-En vs English CIS-Multi | 19.20 | 1.17E-05 | *** | 298 | 143 | 77 | 482 | 18.31 | 1.88E-05 | *** | 278 | 165 | 95 | 1212 |
| Multilingual CIS-Multi vs English CIS-Multi | 18.27 | 1.91E-05 | *** | 246 | 68 | 129 | 557 | 0.07 | 7.97E-01 | 250 | 118 | 123 | 1259 | |
| English CIS-En vs English CIS-Zh | 0.03 | 8.69E-01 | 498 | 72 | 75 | 355 | 0.50 | 4.80E-01 | 288 | 76 | 86 | 1300 | ||
| English CIS-En vs English CIS-Fr | 0.12 | 7.27E-01 | 502 | 68 | 63 | 367 | 0.06 | 8.06E-01 | 291 | 73 | 77 | 1309 | ||
| English CIS-En vs English CIS-Ja | 0.37 | 5.42E-01 | 500 | 70 | 62 | 368 | 0.96 | 3.27E-01 | 283 | 81 | 95 | 1291 | ||
| English CIS-En vs English CIS-Multi | 0.07 | 7.92E-01 | 507 | 63 | 67 | 363 | 0.01 | 9.37E-01 | 284 | 80 | 78 | 1308 | ||
| Multilingual CIS-Multi vs English CIS-Multi | 12.23 | 4.70E-04 | *** | 464 | 63 | 110 | 363 | 0.35 | 5.54E-01 | 266 | 87 | 96 | 1301 | |
| English CIS-En vs English CIS-Zh | 2.06 | 1.51E-01 | 351 | 55 | 40 | 554 | 0.86 | 3.53E-01 | 177 | 52 | 42 | 1479 | ||
| English CIS-En vs English CIS-Fr | 0.01 | 9.23E-01 | 352 | 54 | 54 | 540 | 1.41 | 2.36E-01 | 192 | 37 | 49 | 1472 | ||
| English CIS-En vs English CIS-Ja | 0.34 | 5.62E-01 | 349 | 57 | 50 | 544 | 0.97 | 3.24E-01 | 183 | 46 | 57 | 1464 | ||
| English CIS-En vs English CIS-Multi | 0.00 | 1.00E+00 | 347 | 59 | 58 | 536 | 0.27 | 6.06E-01 | 185 | 44 | 50 | 1471 | ||
| Multilingual CIS-Multi vs English CIS-Multi | 0.03 | 8.55E-01 | 347 | 61 | 58 | 534 | 0.01 | 9.28E-01 | 174 | 63 | 61 | 1452 | ||
| English CIS-En vs English CIS-Zh | 0.01 | 9.42E-01 | 300 | 93 | 95 | 512 | 1.30 | 2.55E-01 | 247 | 83 | 68 | 1352 | ||
| English CIS-En vs English CIS-Fr | 0.20 | 6.57E-01 | 298 | 95 | 88 | 519 | 1.46 | 2.26E-01 | 234 | 96 | 79 | 1341 | ||
| English CIS-En vs English CIS-Ja | 0.08 | 7.80E-01 | 293 | 100 | 105 | 502 | 1.66 | 1.97E-01 | 234 | 96 | 78 | 1342 | ||
| English CIS-En vs English CIS-Multi | 9.66 | 1.88E-03 | ** | 285 | 108 | 66 | 541 | 5.96 | 1.46E-02 | * | 239 | 91 | 60 | 1360 |
| Multilingual CIS-Multi vs English CIS-Multi | 2.31 | 1.28E-01 | 283 | 88 | 68 | 561 | 12.66 | 3.74E-04 | *** | 242 | 103 | 57 | 1348 | |
| English CIS-En vs English CIS-Zh | 0.34 | 5.60E-01 | 646 | 76 | 68 | 210 | 0.02 | 8.78E-01 | 327 | 87 | 84 | 1252 | ||
| English CIS-En vs English CIS-Fr | 0.47 | 4.91E-01 | 650 | 72 | 63 | 215 | 0.56 | 4.55E-01 | 330 | 84 | 95 | 1241 | ||
| English CIS-En vs English CIS-Ja | 1.22 | 2.70E-01 | 634 | 88 | 73 | 205 | 0.35 | 5.52E-01 | 328 | 86 | 95 | 1241 | ||
| English CIS-En vs English CIS-Multi | 0.00 | 1.00E+00 | 662 | 60 | 59 | 219 | 0.02 | 8.77E-01 | 329 | 85 | 82 | 1254 | ||
| Multilingual CIS-Multi vs English CIS-Multi | 6.88 | 8.71E-03 | ** | 614 | 71 | 107 | 208 | 0.68 | 4.09E-01 | 294 | 131 | 117 | 1208 | |
| McNemar’s Test | Low-Resource Languages | High-Resource Languages | ||||||||||||
| XCOPA CIS Mode | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | ||
| M1 vs M2 Comparison | Level | Wrong | M2 Correct | M2 Wrong | Correct | Level | False | M2 Correct | M2 Wrong | Correct | ||||
| English CIS-En vs English CIS-Zh | 199.24 | 3.05E-45 | *** | 1353 | 454 | 116 | 1577 | 8.28 | 4.00E-03 | ** | 349 | 58 | 30 | 2063 |
| English CIS-En vs English CIS-Fr | 214.84 | 1.21E-48 | *** | 1340 | 467 | 113 | 1580 | 3.34 | 6.76E-02 | 349 | 58 | 39 | 2054 | |
| English CIS-En vs English CIS-Ja | 209.37 | 1.89E-47 | *** | 1317 | 490 | 129 | 1564 | 4.90 | 2.69E-02 | * | 341 | 66 | 42 | 2051 |
| English CIS-En vs English CIS-Multi | 278.98 | 1.25E-62 | *** | 1227 | 580 | 133 | 1560 | 13.34 | 2.60E-04 | *** | 330 | 77 | 37 | 2056 |
| Multilingual CIS-Multi vs English CIS-Multi | 0.33 | 5.64E-01 | 1099 | 247 | 261 | 1893 | 0.65 | 4.19E-01 | 285 | 71 | 82 | 2062 | ||
| English CIS-En vs English CIS-Zh | 194.42 | 3.45E-44 | *** | 1133 | 426 | 104 | 1837 | 1.55 | 2.13E-01 | 266 | 53 | 40 | 2141 | |
| English CIS-En vs English CIS-Fr | 39.60 | 3.12E-10 | *** | 1253 | 306 | 168 | 1773 | 0.04 | 8.45E-01 | 268 | 51 | 54 | 2127 | |
| English CIS-En vs English CIS-Ja | 206.75 | 7.04E-47 | *** | 1074 | 485 | 128 | 1813 | 2.44 | 1.18E-01 | 258 | 61 | 44 | 2137 | |
| English CIS-En vs English CIS-Multi | 164.28 | 1.31E-37 | *** | 1077 | 482 | 157 | 1784 | 10.41 | 1.25E-03 | ** | 246 | 73 | 38 | 2143 |
| Multilingual CIS-Multi vs English CIS-Multi | 4.95 | 2.61E-02 | * | 956 | 227 | 278 | 2039 | 0.06 | 8.13E-01 | 206 | 82 | 78 | 2134 | |
| English CIS-En vs English CIS-Zh | 0.13 | 7.15E-01 | 1222 | 97 | 91 | 2090 | 0.01 | 9.07E-01 | 255 | 38 | 36 | 2171 | ||
| English CIS-En vs English CIS-Fr | 0.00 | 9.48E-01 | 1200 | 119 | 119 | 2062 | 1.35 | 2.45E-01 | 268 | 25 | 35 | 2172 | ||
| English CIS-En vs English CIS-Ja | 2.78 | 9.53E-02 | 1184 | 135 | 108 | 2073 | 0.01 | 9.09E-01 | 254 | 39 | 37 | 2170 | ||
| English CIS-En vs English CIS-Multi | 0.14 | 7.09E-01 | 1186 | 133 | 126 | 2055 | 0.05 | 8.15E-01 | 258 | 35 | 38 | 2169 | ||
| Multilingual CIS-Multi vs English CIS-Multi | 2.96 | 8.55E-02 | 1099 | 178 | 213 | 2010 | 9.60 | 1.95E-03 | ** | 210 | 49 | 86 | 2155 | |
| English CIS-En vs English CIS-Zh | 4.03 | 4.48E-02 | * | 1092 | 161 | 126 | 2121 | 2.53 | 1.12E-01 | 216 | 35 | 22 | 2227 | |
| English CIS-En vs English CIS-Fr | 0.01 | 9.10E-01 | 1098 | 155 | 158 | 2089 | 1.84 | 1.75E-01 | 198 | 53 | 39 | 2210 | ||
| English CIS-En vs English CIS-Ja | 0.00 | 9.57E-01 | 1078 | 175 | 175 | 2072 | 0.86 | 3.53E-01 | 199 | 52 | 42 | 2207 | ||
| English CIS-En vs English CIS-Multi | 2.06 | 1.51E-01 | 1034 | 219 | 189 | 2058 | 3.14 | 7.63E-02 | 196 | 55 | 37 | 2212 | ||
| Multilingual CIS-Multi vs English CIS-Multi | 2.08 | 1.49E-01 | 1008 | 247 | 215 | 2030 | 0.92 | 3.38E-01 | 161 | 60 | 72 | 2207 | ||
| English CIS-En vs English CIS-Zh | 0.05 | 8.24E-01 | 1201 | 164 | 159 | 1976 | 1.87 | 1.72E-01 | 303 | 60 | 45 | 2092 | ||
| English CIS-En vs English CIS-Fr | 0.13 | 7.15E-01 | 1177 | 188 | 180 | 1955 | 1.98 | 1.59E-01 | 321 | 42 | 57 | 2080 | ||
| English CIS-En vs English CIS-Ja | 1.78 | 1.82E-01 | 1146 | 219 | 191 | 1944 | 0.08 | 7.84E-01 | 305 | 58 | 62 | 2075 | ||
| English CIS-En vs English CIS-Multi | 4.27 | 3.88E-02 | * | 1137 | 228 | 185 | 1950 | 10.54 | 1.17E-03 | ** | 283 | 80 | 43 | 2094 |
| Multilingual CIS-Multi vs English CIS-Multi | 0.12 | 7.24E-01 | 1069 | 262 | 253 | 1916 | 0.88 | 3.47E-01 | 238 | 75 | 88 | 2099 | ||
| English CIS-En vs English CIS-Zh | 70.71 | 4.14E-17 | *** | 2503 | 185 | 54 | 758 | 36.62 | 1.43E-09 | *** | 310 | 100 | 30 | 2060 |
| English CIS-En vs English CIS-Fr | 124.96 | 5.19E-29 | *** | 2444 | 244 | 51 | 761 | 15.24 | 9.45E-05 | *** | 324 | 86 | 41 | 2049 |
| English CIS-En vs English CIS-Ja | 275.84 | 6.05E-62 | *** | 2301 | 387 | 42 | 770 | 34.30 | 4.73E-09 | *** | 313 | 97 | 30 | 2060 |
| English CIS-En vs English CIS-Multi | 298.12 | 8.46E-67 | *** | 2285 | 403 | 39 | 773 | 57.51 | 3.37E-14 | *** | 292 | 118 | 26 | 2064 |
| Multilingual CIS-Multi vs English CIS-Multi | 294.90 | 4.26E-66 | *** | 1795 | 98 | 529 | 1078 | 9.14 | 2.50E-03 | ** | 233 | 49 | 85 | 2133 |
| McNemar’s Test | Low-Resource Languages | High-Resource Languages | ||||||||||||
| XL-WiC CIS Mode | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | ||
| M1 vs M2 Comparison | Level | Wrong | M2 Correct | M2 Wrong | Correct | Level | False | M2 Correct | M2 Wrong | Correct | ||||
| English CIS-En vs English CIS-Zh | 3.31 | 6.90E-02 | 583 | 92 | 68 | 817 | 1.69 | 1.94E-01 | 1251 | 183 | 158 | 1918 | ||
| English CIS-En vs English CIS-Fr | 0.81 | 3.69E-01 | 594 | 81 | 69 | 816 | 1.16 | 2.81E-01 | 1269 | 165 | 145 | 1931 | ||
| English CIS-En vs English CIS-Ja | 2.25 | 1.34E-01 | 566 | 109 | 87 | 798 | 2.96 | 8.55E-02 | 1221 | 213 | 178 | 1898 | ||
| English CIS-En vs English CIS-Multi | 1.29 | 2.57E-01 | 567 | 108 | 91 | 794 | 0.00 | 1.00E+00 | 1226 | 208 | 207 | 1869 | ||
| Multilingual CIS-Multi vs English CIS-Multi | 0.13 | 7.20E-01 | 464 | 186 | 194 | 716 | 4.03 | 4.46E-02 | * | 1044 | 334 | 389 | 1743 | |
| English CIS-En vs English CIS-Zh | 34.03 | 5.43E-09 | *** | 598 | 215 | 109 | 638 | 67.56 | 2.04E-16 | *** | 1234 | 414 | 208 | 1654 |
| English CIS-En vs English CIS-Fr | 16.94 | 3.86E-05 | *** | 604 | 209 | 132 | 615 | 72.71 | 1.50E-17 | *** | 1298 | 350 | 157 | 1705 |
| English CIS-En vs English CIS-Ja | 47.64 | 5.13E-12 | *** | 586 | 227 | 101 | 646 | 52.87 | 3.56E-13 | *** | 1207 | 441 | 249 | 1613 |
| English CIS-En vs English CIS-Multi | 45.23 | 1.75E-11 | *** | 574 | 239 | 112 | 635 | 53.72 | 2.31E-13 | *** | 1171 | 477 | 275 | 1587 |
| Multilingual CIS-Multi vs English CIS-Multi | 0.97 | 3.25E-01 | 509 | 158 | 177 | 716 | 0.83 | 3.63E-01 | 992 | 426 | 454 | 1638 | ||
| English CIS-En vs English CIS-Zh | 26.89 | 2.16E-07 | *** | 566 | 200 | 108 | 686 | 200.56 | 1.58E-45 | *** | 1772 | 391 | 82 | 1265 |
| English CIS-En vs English CIS-Fr | 8.28 | 4.00E-03 | ** | 638 | 128 | 85 | 709 | 32.76 | 1.04E-08 | *** | 1984 | 179 | 85 | 1262 |
| English CIS-En vs English CIS-Ja | 19.38 | 1.07E-05 | *** | 551 | 215 | 132 | 662 | 242.30 | 1.24E-54 | *** | 1695 | 468 | 97 | 1250 |
| English CIS-En vs English CIS-Multi | 14.08 | 1.75E-04 | *** | 598 | 168 | 105 | 689 | 61.98 | 3.46E-15 | *** | 1906 | 257 | 106 | 1241 |
| Multilingual CIS-Multi vs English CIS-Multi | 0.34 | 5.61E-01 | 446 | 243 | 257 | 614 | 392.40 | 2.48E-87 | *** | 1089 | 245 | 923 | 1253 | |
| English CIS-En vs English CIS-Zh | 0.03 | 8.64E-01 | 569 | 70 | 67 | 854 | 1.77 | 1.83E-01 | 1060 | 148 | 125 | 2177 | ||
| English CIS-En vs English CIS-Fr | 0.01 | 9.36E-01 | 561 | 78 | 78 | 843 | 0.01 | 9.07E-01 | 1063 | 145 | 148 | 2154 | ||
| English CIS-En vs English CIS-Ja | 0.10 | 7.48E-01 | 559 | 80 | 75 | 846 | 2.30 | 1.29E-01 | 1023 | 185 | 156 | 2146 | ||
| English CIS-En vs English CIS-Multi | 0.09 | 7.70E-01 | 543 | 96 | 91 | 830 | 2.01 | 1.57E-01 | 1026 | 182 | 155 | 2147 | ||
| Multilingual CIS-Multi vs English CIS-Multi | 4.75 | 2.92E-02 | * | 494 | 180 | 140 | 746 | 0.24 | 6.27E-01 | 883 | 311 | 298 | 2018 | |
| English CIS-En vs English CIS-Zh | 0.71 | 4.01E-01 | 757 | 29 | 22 | 752 | 16.69 | 4.39E-05 | *** | 1373 | 173 | 104 | 1860 | |
| English CIS-En vs English CIS-Fr | 4.17 | 4.11E-02 | * | 755 | 31 | 16 | 758 | 16.67 | 4.45E-05 | *** | 1354 | 192 | 119 | 1845 |
| English CIS-En vs English CIS-Ja | 0.20 | 6.51E-01 | 762 | 24 | 20 | 754 | 26.40 | 2.77E-07 | *** | 1351 | 195 | 105 | 1859 | |
| English CIS-En vs English CIS-Multi | 0.93 | 3.36E-01 | 747 | 39 | 30 | 744 | 23.18 | 1.47E-06 | *** | 1343 | 203 | 116 | 1848 | |
| Multilingual CIS-Multi vs English CIS-Multi | 10.28 | 1.35E-03 | ** | 683 | 54 | 94 | 729 | 9.39 | 2.18E-03 | ** | 1130 | 254 | 329 | 1797 |
| English CIS-En vs English CIS-Zh | 7.64 | 5.72E-03 | ** | 521 | 137 | 94 | 808 | 102.12 | 5.24E-24 | *** | 1402 | 277 | 84 | 1747 |
| English CIS-En vs English CIS-Fr | 10.90 | 9.60E-04 | *** | 480 | 178 | 120 | 782 | 146.30 | 1.12E-33 | *** | 1292 | 387 | 115 | 1716 |
| English CIS-En vs English CIS-Ja | 9.03 | 2.66E-03 | ** | 500 | 158 | 108 | 794 | 94.02 | 3.12E-22 | *** | 1404 | 275 | 89 | 1742 |
| English CIS-En vs English CIS-Multi | 14.59 | 1.34E-04 | *** | 465 | 193 | 124 | 778 | 143.32 | 5.01E-33 | *** | 1301 | 378 | 112 | 1719 |
| Multilingual CIS-Multi vs English CIS-Multi | 0.80 | 3.72E-01 | 417 | 190 | 172 | 781 | 43.63 | 3.96E-11 | *** | 965 | 270 | 448 | 1827 | |

### v3

@@
 Among several prompting strategies aiming at bridging the gap, multilingual in-context learning (ICL; Shi et al., 2023) has been particularly effective when demonstration in target languages is unavailable.
-However, there lacks a systematic understanding when and why it works well.
+However, there lacks a systematic understanding of when and why it works well.
 In this work, we systematically analyze multilingual ICL, using demonstrations in HRLs to enhance cross-lingual transfer.
 We show that demonstrations in mixed HRLs consistently outperform English-only ones across the board, particularly for tasks written in LRLs.
-Surprisingly, our ablation study show that the presence of irrelevant non-English sentences in the prompt yields measurable gains, suggesting the effectiveness of multilingual exposure itself.
+Surprisingly, our ablation study shows that the presence of irrelevant non-English sentences in the prompt yields measurable gains, suggesting the effectiveness of multilingual exposure itself.
 Our results highlight the potential of strategically leveraging multilingual resources to bridge the performance gap for underrepresented languages.
 Blessing of Multilinguality: A Systematic Analysis of Multilingual In-Context Learning
-Yilei Tu Andrew Xue Freda Shi David R. Cheriton School of Computer Science, University of Waterloo 200 University Ave W, Waterloo, ON N2L 5Z5, Canada {yileitu,a8xue,fhs}@uwaterloo.ca
+Yilei Tu1 Andrew Xue1 Freda Shi1,2,3 1David R. Cheriton School of Computer Science, University of Waterloo 2Vector Institute 3Canada CIFAR AI Chair yileitu.tt@gmail.com, fhs@uwaterloo.ca
 1 Introduction
@@
 Both strategies have critical shortcomings: (i) may suffer from information loss in translation due to nuanced language gap (Zhang et al., 2023; Poelman and de Lhoneux, 2024) or the unavailability of high-quality translation systems for extremely low-resource languages (LRLs), whereas (ii) may become infeasible due to data scarcity in LRLs.
-As alternatives, when presenting LLMs with the problems in the target language, English demonstrations (Fig.
-1(a)) lead to poor performance on LRLs, whereas demonstrations in multiple high-resource languages (HRLs; Fig.
-1(b)) can be more effective, even when there are little alphabetical overlap between the demonstration and target languages (Shi et al., 2023); however, the underlying reasons on why it works remain unclear.
+As alternatives, when presenting LLMs with problems in the target language, English demonstrations (Fig.˜1(a)) lead to poor performance on LRLs, whereas demonstrations in multiple high-resource languages (HRLs; Fig.˜1(b)) can be more effective, even when there is little alphabetical overlap between the demonstration and target languages (Shi et al., 2023); however, the underlying reasons on why it works remain unclear.
 In this work, we systematically analyze multilingual ICL through a set of controlled experiments.
-Each test question is paired with a set of semantically equivalent demonstrations, while the demonstration languages vary according to different ICL modes (Tab.
-1).
-We compare the performance differences across four ICL modes (§§ 4.1 and 4.2): English, individual-HRL, multilingual (i.e., mixed HRLs), and in-language demonstrations (Fig.
-2).
-To disentangle the impact of demonstration language from other confounding factors (e.g., interactions between demonstration languages and in-domain demonstrations), we conduct additional control experiments by adding irrelevant sentences in various languages into English-only in-domain demonstrations (§ 4.3).
+Each test question is paired with a set of semantically equivalent demonstrations, while the demonstration languages vary according to different ICL modes (Tab.˜1).
+We compare the performance differences across four ICL modes (§§˜4.1 and 4.2): English, individual-HRL, multilingual (i.e., mixed HRLs), and in-language demonstrations (Fig.˜2).
+To disentangle the impact of demonstration language from other confounding factors (e.g., interactions between demonstration languages and in-domain demonstrations), we conduct additional control experiments by adding irrelevant sentences in various languages into English-only in-domain demonstrations (§˜4.3).
 We find that
-In-context demonstrations in HRLs, especially in languages with non-Latin writing system such as Chinese and Japanese, can more effectively transfer knowledge compared to English ICL mode, leading to performance improvement on answering questions in all languages, especially in LRLs.
+In-context demonstrations in HRLs, especially in languages with non-Latin writing systems such as Chinese and Japanese, can more effectively transfer knowledge compared to English ICL mode, leading to performance improvement in answering questions in all languages, especially in LRLs.
 This finding is generalizable enough across different domains and various LLMs.
-Demonstrations in mixed HRLs is more robust and effective compared to that in a single HRL, in terms of average accuracy boosting on different tasks. This strategy is favored.
+Demonstrations in mixed HRLs are more robust and effective compared to those in a single HRL, in terms of average accuracy boosting on different tasks. This strategy is favored.
 Surprisingly, simply introducing another non-English language (not necessarily in the demonstration) in the prompt could lead to performance improvement, albeit the improvement is not as significant as the aforementioned strategies.
 2 Background: Multilingual ICL
-In this section, we review the basic approaches of ICL with instruction-tuned LLMs (§ 2.1) and introduce the multilingual prompting modes (§ 2.2) that we evaluate in this work.
+In this section, we review the basic approaches of ICL with instruction-tuned LLMs (§˜2.1) and introduce the multilingual prompting modes (§˜2.2) that we evaluate in this work.
 2.1 ICL for Instruction-Tuned LLM
@@
 In such cases, the verbalizer is an identity.
-For answers requiring reasoning steps, is CoT in English, as there has been strong evidence that MLLMs performs better when generating English (Shi et al., 2023; Qin et al., 2023; Huang et al., 2023, inter alia).
+For answers requiring reasoning steps, is CoT in English, as there has been strong evidence that MLLMs perform better when generating English (Shi et al., 2023; Qin et al., 2023; Huang et al., 2023, inter alia).
 For the same reasons, is always presented in English as well.
 Following Ahuja et al. (2023) and Shi et al. (2023), we evaluate MLLMs via several different prompting strategies (ICL modes) in this work:
-The English mode. The demonstrations are always in English (Figs. 1(a) and 2a).
-The Monolingual mode(s). The demonstrations are always presented in a single non-English high-resource language such as Chinese (Fig. 2b).
+The English mode. The demonstrations are always in English (Figs.˜1(a) and 2a).
+The Monolingual mode(s). The demonstrations are always presented in a single non-English high-resource language such as Chinese (Fig.˜2b).
 The Multilingual mode.
-From a predefined list of high-resources languages , languages are randomly selected, which, together with the sampled indices, determine the contents and languages of the demonstrations (Figs.
-1(b) and 2c).
-The Native mode. The demonstrations are in the same language as the test question (Fig. 2d).
+From a predefined list of high-resources languages , languages are randomly selected, which, together with the sampled indices, determine the contents and languages of the demonstrations (Figs.˜1(b) and 2c).
+The Native mode. The demonstrations are in the same language as the test question (Fig.˜2d).
 3 Experiment Setups
@@
 For additional references, we evaluate OpenAI closed-sourced commercial models, including GPT3.5-turbo (OpenAI, 2022) and GPT4o-mini (OpenAI, 2024).
-Detailed model cards can be found in § A.1.
+Detailed model cards can be found in §˜A.1.
 | Dataset | Domain | Expected | #Languages | Language Split Size | En Avg Word | Parallel |
 | Datapoint Example | Output | (#HRL #LRL) | — Training/Test | Count | ||
-[…]
-| See Fig. 1 for examples. | ||||||
-[…]
+| MGSM | Mathematical Reasoning | Numerals | ||||
+| See Fig.˜1 for examples. | ||||||
+| XCOPA | Commonsense Reasoning | “1” / “2” | ||||
 | Premise: The man turned on the faucet. What happened as a RESULT? | ||||||
 | Hypothesis 1: The toilet filled with water. Hypothesis 2: Water flowed from the spout. Answer: 2. | ||||||
-[…]
+| XL-WiC | Word Disambiguation | “Yes” / “No” | ||||
 | Sentence 1: What did you *get* at the toy store? Sentence 2: She didn’t *get* his name when they met the first time. | ||||||
 | Question: Is the word “get” (marked with *) used in the same way in both sentences above? Answer: No. | ||||||
+| XQuAD | Extractive QA | Text | ||||
+| Passage: The Panthers defense gave up just 308 points, ranking sixth in the league, | ||||||
+| Question: How many points did the Panthers defense surrender? Answer: 308. | ||||||
 Datasets.
-We evaluate the models using multilingual benchmarks from three distinct domains: (1) MGSM Shi et al.
-(2023), a benchmark of 250 grade-school math problems sampled from the English GSM8K (Cobbe et al., 2021) and translated into 10 additional languages by expert native speakers.
+We evaluate the models using multilingual benchmarks from distinct domains: (1) MGSM Shi et al.
+(2023), a benchmark of grade-school math problems sampled from the English GSM8K (Cobbe et al., 2021) and translated into 10 additional languages by expert native speakers.
 (2) XCOPA Ponti et al.
-(2020), a commonsense reasoning benchmark that extends the COPA dataset (Roemmele et al., 2011) to additional languages.111In this work, we merge the English COPA and XCOPA datasets together, which we still refer to as XCOPA.
+(2020), a commonsense reasoning benchmark that extends the COPA dataset in English (Roemmele et al., 2011) to additional languages.
 (3) XL-WiC Raganato et al.
 (2020), a cross-lingual word-in-context understanding dataset spanning languages, where models are expected to tell whether a polysemous word retains the same meaning in two contexts.
-MGSM and XCOPA are parallel where each corresponding datapoint across different language splits contains semantically equivalent content, allowing us to minimize semantic confounders in our experimental design.
+(4) XQuAD Artetxe et al.
+(2020) is designed to evaluate cross-lingual question answering performance, based on the English SQuAD Rajpurkar et al.
+(2016) dataset and professionally translated into 10 languages.
+MGSM, XCOPA and XQuAD are parallel where each corresponding datapoint across different language splits contains semantically equivalent content, allowing us to minimize semantic confounders in our experimental design.
 XL-WiC is language-specific and translation-variant thus naturally non-parallel.
-Dataset properties and examples are in Tabs.
-1 and 6.
-Details of data curation are in § A.2.
+Dataset properties and examples are in Tabs.˜1 and 6.
+Details of data curation are in §˜A.2.
 Languages.
-Languages with large-scale digitized data resources on web are known as high-resource languages (HRLs; Bender, 2019), which are exemplified by English, Spanish, Chinese, among others.
+Languages with large-scale digitized data resources on the web are known as high-resource languages (HRLs; Bender, 2019), which are exemplified by English, Spanish, and Chinese, among others.
 In contrast, low-resource languages (LRL) have scarce accessible data Costa-jussà et al.
@@
 (2023), and classify languages out of the HRL list as LRLs.
-Details of preset HRL list can be found in § A.3.
+Details of the preset HRL list can be found in §˜A.3.
 Prompts.
@@
 (2023), we use examples for demonstration for any test question.
-For each multilingual dataset (Tab.
-1), we first sample index lists of length all at once, where the index range is We allocate the -th index list to the set of test questions for the same index across all language splits.
-The training-set index list, together with the specified languages,222 For the Multilingual mode, we apply the same sampling procedure to generate HRL code lists of length , drawing from the available HRLs in the dataset.
-jointly determines the content and language of the demonstration for each testing example (Fig.
-2).
+For each multilingual dataset (Tab.˜1), we first sample index lists of length all at once, where the index range is We allocate the -th index list to the set of test questions for the same index across all language splits.
+The training-set index list, together with the specified languages,111For the Multilingual mode, we apply the same sampling procedure to generate HRL code lists of length , drawing from the available HRLs in the dataset.
+jointly determines the content and language of the demonstration for each testing example (Fig.˜2).
 This approach both ensures linguistic diversity for multilingual prompting and, whenever applicable, mitigates confounding factors that come with semantic inconsistency across examples.
-All interface words (see Tab.
-1) are in the same language as the examples rather than in English.
+All interface words (see Tab.˜1) are in the same language as the examples rather than in English.
 Inference and metrics.
 Throughout this work, we use greedy decoding for inference, selecting the token with the highest probability at each step.
-For MGSM in need of CoT, we set the maximum token length to 500; for XCOPA and XL-WiC, we set it to 10, as we expect the answers to be short.
-We use exact match accuracy as our evaluation metric: for MGSM, we extract the last numeral in the response.
-For XCOPA and XL-WiC, we extract label (expected output) in the response (Tab.
-1).
+For MGSM in need of CoT, we set the maximum token length to ; for XCOPA and XL-WiC, we set it to 10, as we expect the answers to be short; for XQuAD, it is .
+We use exact match accuracy as our evaluation metric: for MGSM, we extract the last numeral in the response; for XCOPA and XL-WiC, we extract the label (expected output) in the response (Tab.˜1); for XQuAD, a response of an MLLM is considered correct if it contains the gold answer as a substring after normalization.
 4 ICL Mode Evaluation
@@
 Results.
-We first compare the -shot performance with English, Multilingual and Native ICL modes on open-sourced MLLMs and commercial OpenAI models (Fig.
-3), with Tab.
-2 presenting the detailed performance of three selected MLLMs across various LRLs.
+We first compare the -shot performance with English, Multilingual and Native ICL modes on open-sourced MLLMs and commercial OpenAI models (Fig.˜3), with Tab.˜2 presenting the detailed performance of three selected MLLMs across various LRLs.
 Overall, Multilingual mode outperforms English mode, both for individual LRLs and on average.
-In out of cases (Fig.
-3), Multilingual mode achieves higher accuracy than English one.
+In out of cases (Fig.˜3), Multilingual mode achieves higher accuracy than English one.
 This phenomenon is evident even for GPT4o-mini, one of the currently strongest LLMs Chiang et al.
-(2024), and for HRLs as well (see Tabs.
-11, 12 and 13 in § B.1 for HRL accuracies).
+(2024), and for HRLs as well (see Tabs.˜11, 12 and 13 in §˜B.1 for HRL accuracies).
 Extending the results of Shi et al.
@@
 Outstanding performance of the Native mode and the practical unfeasibility.
-Admittedly, in out of comparisons, Native mode outperforms Multilingual mode (Fig.
-3), which aligns with the machine-learning intuition that in-domain data, in terms of both genre and language, are more promising for model performance (Liu et al., 2024).
+Admittedly, in out of comparisons, Native mode outperforms Multilingual mode (Fig.˜3), which aligns with the machine-learning intuition that in-domain data, in terms of both genre and language, are more promising for model performance (Liu et al., 2024).
 However, domain-specific datasets for LRLs are often difficult to obtain due to the scarcity of native speakers or professional translators (Costa-jussà et al., 2022; NLLB-Team, 2024); therefore, in practice, it is usually challenging to provide high-quality demonstrations in the same language and domain as the test question.
-In contrast, annotations makes the HRL-Multilingual mode more feasible in many scenarios.
+In contrast, annotations make the HRL-Multilingual mode more feasible in many scenarios.
 Hypothesis Tests.
 To verify whether the improvement is statistically significant, we conduct McNemar’s test McNemar (1947)—the null hypothesis means no significant accuracy difference between the baseline (English) and the compared mode.
-Let denote the number of cases where the baseline is correct while the compared mode is incorrect, denotte the number of cases where the baseline mode is incorrect while the compared mode is correct.
+Let denote the number of cases where the baseline is correct while the compared mode is incorrect, and denote the number of cases where the baseline mode is erroneous while the compared mode is correct.
 We calculate the corrected version Edwards (1948) of the McNemar’s statistic:
@@
 Significant -test results provide strong evidence to reject the null hypothesis of no accuracy improvement.
-Our results results (Tab.
-2 and Tabs.
-14, 15 and 16 in § B.1) indicate that both Multilingual and Native modes significantly outperform the English mode.
+Our results results (Tab.˜2 and Tabs.˜15, 16, 17 and 18 in §˜B.1) indicate that both Multilingual and Native modes significantly outperform the English mode.
 4.2 Ablation Study: Non-English Monolingual Prompts Are Effective
 With the success of the Multilingual mode, we investigate whether the improvement comes from the introduction of multiple languages or simply from a single non-English HRL.
-Specifically, we compare several Monolingual modes including Chinese (for all three datasets), French, Japanese (both for MGSM and XL-WiC) and Italian (for XCOPA).
-We select French and Italian because they are both European languages and thus share considerable subword overlap with English; in contrast, Chinese and Japanese exhibit little subword overlap with Latin-script languages, but there is a substantial overlap between them two due to their shared use of Chinese characters (or kanji).
+Specifically, we compare several Monolingual modes including Chinese (for all datasets), French, Japanese (both for MGSM and XL-WiC), Italian (for XCOPA) and German (for XQuAD).
+We select French, Italian and German because they are European languages and thus share considerable subword overlap with English; in contrast, Chinese and Japanese exhibit little subword overlap with Latin-script languages, but there is a substantial overlap between the two due to their shared use of Chinese characters (or Hanzi, Kanji).
 This language selection allows us to analyze simultaneously the impact of the writing system (or subword overlap) on ICL performance.
 Results.
-All HRL-Monolingual modes outperform English in a considerable number of comparisons (Fig.
-4).
-This finding also holds for HRL evaluations (see Tabs.
-11, 12 and 13 in § B.1 for raw accuracies of each language).
-Among all Monolingual modes, Chinese performs the best, matching Multilingual mode with 17 out of 21 comparisons when accuracy surpasses that of English.
+All HRL-Monolingual modes outperform English in a considerable number of comparisons (Fig.˜4).
+This finding also holds for HRL evaluations (see Tabs.˜11, 12, 13 and 14 in §˜B.1 for raw accuracies of each language).
+Among all Monolingual modes, Chinese performs the best, with out of comparisons when accuracy surpasses that of English.
 Japanese also frequently outperforms English.
@@
 (2021) that non-English languages are more effective than English in pretraining and fine-tuning based cross-lingual transfer, our results suggest that non-English languages, particularly those with non-Latin scripts, may be more effective under the prompting scheme as well.
-However, the Multilingual mode exhibits stronger robustness, outperforming Chinese in 14 out of 21 LRL cases.
-Same trends applies to HRLs.
-The results of the hypothesis test further confirm the robustness: for both LRL and HRL splits, the Multilingual mode exhibits the highest number of significant results relative to other Monolingual modes (Tabs.
-14, 15 and 16 in § B.1).
+However, the Multilingual mode exhibits stronger robustness, outperforming Chinese in out of LRL cases.
+The same trend applies to HRLs.
+The results of the hypothesis test further confirm the robustness: for both LRL and HRL splits, the Multilingual mode exhibits the highest number of significant results relative to other Monolingual modes (Tabs.˜15, 16, 17 and 18 in §˜B.1).
 Intuitively, we hypothesize that Multilingual mode functions like an “average” of individual HRL-Monolingual modes, making it most robust, and thus it outperforms English most frequently and achieves the highest overall average accuracy.
@@
 This further explains why Multilingual could achieve performance comparable to Native.
-See App.
-C for details.
+See App.˜C for details.
 4.3 Ablation Study: Merely Introducing New Language(s) Enhances ICL Performance
@@
 To distinguish these two setups, we prepend a context-irrelevant sentence (CIS) before each ICL demonstration, which is unrelated to the current domain and can be in any language.
-Analogous to the ICL modes introduced in § 2.2, CIS resembles the construction of those modes with the same sampling strategy (§ 3).
-For example, based on the English mode, we could prepend a set of multilingual CIS, which augments Fig.
-2a into Fig.
-5.
+Analogous to the ICL modes introduced in §˜2.2, CIS resembles the construction of those modes with the same sampling strategy (§˜3).
+For example, based on the English mode, we could prepend a set of multilingual CIS, which augments Fig.˜2a into Fig.˜5.
 We denote such setting by “English CIS-Multi”.
@@
 (2022) as the source of our CIS, which provides parallel Wikipedia sentences in multiple languages, a fairly distant genre to all evaluation datasets.
-We filter sentence-parallel sets with English word counts ranging from 10 to 15 as our sampling pool , a small range compared to target datasets (Tab.
-1), to mitigate the risk of introducing too much noise.
-Details of FLORES-101 can be found at § A.2.
+We filter sentence-parallel sets with English word counts ranging from 10 to 15 as our sampling pool , a small range compared to target datasets (Tab.˜1), to mitigate the risk of introducing too much noise.
+Details of FLORES-101 can be found at §˜A.2.
 Results and analysis.
-We perform hypothesis testing as in Tab.
-2 to compare different CIS settings with English CIS-En (Tab.
-3)—English CIS-En exhibit a small drop compared to English only, indicating that our filtration effectively controls the negative impact of noise to a tolerable level.
+We perform hypothesis testing as in Tab.˜2 to compare different CIS settings with English CIS-En (Tab.˜3)—English CIS-En exhibit a small drop compared to English only, indicating that our filtration effectively controls the negative impact of noise to a tolerable level.
 We find that introducing a single non-English language generally improves ICL performance in most cases ; however, the improvement is only statistically significant in cases.
 This reveals that simply introducing a language can lead to a modest improvement in MLLM performance, but it is more pronounced when in-topic demonstrations in another language (Monolingual modes) are incorporated.
-More hypothesis test results for both LRL and HRL splits can be found in Tabs.
-20, 21 and 22 in § B.2.
-Introducing multiple languages (CIS-Multi, Fig.
-5) is slightly more promising than CIS in a single language, with of cases showing improvement, of which are significant.
-We further conduct experiments by prepending multilingual CIS to the Multilingual mode (Multilingual CIS-Multi, Tab.
-3).
-The performance of Multilingual CIS-Multi is not significantly lower than that of English CIS-Multi; in more than half cases, it significantly outperforms English CIS-Multi.
+More hypothesis test results for both LRL and HRL splits can be found in Tabs.˜23, 24 and 25 in §˜B.2.
+Introducing multiple languages (CIS-Multi, Fig.˜5) is slightly more promising than CIS in a single language, with of cases showing improvement, of which are significant.
+We further conduct experiments by prepending multilingual CIS to the Multilingual mode (Multilingual CIS-Multi, Tab.˜3).
+The performance of Multilingual CIS-Multi is not significantly lower than that of English CIS-Multi; in cases, it significantly outperforms English CIS-Multi.
 This concludes that the significant improvement from English mode to Multilingual mode is attributed to the incorporation of multiple languages with in-topic demonstrations.
-| LRL AvgΔ↑↓ | MGSM | XL-WiC | XCOPA |
-| English | 64.20 | 57.37 | 46.23 |
-| English CIS-En | 61.00 | 56.73 | 48.37 |
-| English CIS-Fr | |||
-| English CIS-Ja | |||
-| English CIS-Zh | |||
-| English CIS-Multi | |||
-| Multilingual CIS-Multi | |||
-| English | 57.10 | 44.42 | 55.91 |
-| English CIS-En | 55.90 | 47.88 | 55.46 |
-| English CIS-Fr | |||
-| English CIS-Ja | |||
-| English CIS-Zh | |||
-| English CIS-Multi | |||
-| Multilingual CIS-Multi | |||
-| English | 43.70 | 48.46 | 62.29 |
-| English CIS-En | 43.00 | 50.90 | 62.31 |
-| English CIS-Fr | |||
-| English CIS-Ja | |||
-| English CIS-Zh | |||
-| English CIS-Multi | |||
-| Multilingual CIS-Multi | |||
-| English | 59.40 | 57.82 | 64.69 |
-| English CIS-En | 59.40 | 59.04 | 64.20 |
-| English CIS-Fr | |||
-| English CIS-Ja | |||
-| English CIS-Zh | |||
-| English CIS-Multi | |||
-| Multilingual CIS-Multi | |||
-| English | 57.00 | 51.54 | 62.26 |
-| English CIS-En | 60.70 | 49.62 | 61.00 |
-| English CIS-Fr | |||
-| English CIS-Ja | |||
-| English CIS-Zh | |||
-| English CIS-Multi | |||
-| Multilingual CIS-Multi | |||
-| English | 29.40 | 58.78 | 26.54 |
-| English CIS-En | 27.80 | 57.82 | 23.20 |
-| English CIS-Fr | |||
-| English CIS-Ja | |||
-| English CIS-Zh | |||
-| English CIS-Multi | |||
-| Multilingual CIS-Multi | |||
+| LRL | MGSM | XL-WiC | XCOPA | XQuAD |
+| English | 64.20 | 57.37 | 46.23 | 72.02 |
+| CIS-En | 61.00 | 56.73 | 48.37 | 71.43 |
+| CIS-Fr | ||||
+| CIS-Ja | ||||
+| CIS-Zh | ||||
+CIS- |
+| English | 57.10 | 44.42 | 55.91 | 68.25 |
+| CIS-En | 55.90 | 47.88 | 55.46 | 68.45 |
+| CIS-Fr | ||||
+| CIS-Ja | ||||
+| CIS-Zh | ||||
+CIS- |
+| English | 43.70 | 48.46 | 62.29 | 53.40 |
+| CIS-En | 43.00 | 50.90 | 62.31 | 58.55 |
+| CIS-Fr | ||||
+| CIS-Ja | ||||
+| CIS-Zh | ||||
+CIS- |
+| English | 59.40 | 57.82 | 64.69 | 68.98 |
+| CIS-En | 59.40 | 59.04 | 64.20 | 69.05 |
+| CIS-Fr | ||||
+| CIS-Ja | ||||
+| CIS-Zh | ||||
+CIS- |
+| English | 57.00 | 51.54 | 62.26 | 60.93 |
+| CIS-En | 60.70 | 49.62 | 61.00 | 56.48 |
+| CIS-Fr | ||||
+| CIS-Ja | ||||
+| CIS-Zh | ||||
+CIS- |
+| English | 29.40 | 58.78 | 26.54 | 70.48 |
+| CIS-En | 27.80 | 57.82 | 23.20 | 70.30 |
+| CIS-Fr | ||||
+| CIS-Ja | ||||
+| CIS-Zh | ||||
+CIS- |
 4.4 Translation-Based Performance
@@
 We test two translation strategies for baseline comparison.
-(1) Transl-LangEn: Translating test questions in other languages in the English ICL mode (Fig.
-2a) into English.
-(2) Transl-EnLang: Translating demonstrations in English ICL mode (Fig.
-2a) into the source language of the current test question, which mirrors the Native mode (Fig.
-2d).
-We use the Google Cloud Translation API for translation.333https://cloud.google.com/translate/
+(1) Transl-LangEn: Translating test questions in other languages in the English ICL mode (Fig.˜2a) into English.
+(2) Transl-EnLang: Translating demonstrations in English ICL mode (Fig.˜2a) into the source language of the current test question, which mirrors the Native mode (Fig.˜2d).
+We use the Google Cloud Translation API for translation.222https://cloud.google.com/translate/
 | Avg Acc | MGSM | XCOPA | ||
@@
 Analysis.
-For LRLs, Transl-LangEn sometimes outperform the Native mode, while Transl-EnLang performs comparably to the Multilingual mode, but falls short of the Native mode performance (Tab.
-4).
+For LRLs, Transl-LangEn sometimes outperform the Native mode, while Transl-EnLang performs comparably to the Multilingual mode, but falls short of the Native mode performance (Tab.˜4).
 These results resonate with the phenomena that the translation quality of LRLEn is generally higher than that of the reversed direction Fan et al.
@@
 (2023).
+Concurrent work Gao et al.
+(2025) shows that multilingual prompting has improved the general reasoning capabilities of multilingual LLMs.
 Multilingual ICL.
-For languages of templates, demonstration and sample question in native language is conventionally inserted into a predefined English template Lin et al.
+For languages of templates, demonstrations and sample questions in native languages are conventionally inserted into a predefined English template Lin et al.
 (2022); Fu et al.
@@
 This paper treats multilingual LLMs as black-box models, drawing the findings and conclusions based solely on their input-output behavior.
-Hence we have not interpreted the internal mechanism how multilinguality could affect MLLM’s “thinking” process and its manifested performance.
-We have briefly touched the impact of demonstrations in different languages on the MLLM performance.
+Hence, we have not interpreted the internal mechanism of how multilingualism could affect MLLM’s “thinking” process and its manifested performance.
+We have briefly touched on the impact of demonstrations in different languages on the MLLM performance.
 However, we do not conduct a thorough empirical analysis to identify which specific linguistic characteristics (e.g., writing systems, grammatical structures, or linguistic relatedness) contribute to the observed performance differences.
-Reproducibility
-Our code is available at https://github.com/yileitu/multilingual_icl.
 Appendix A Experiment Setup
@@
 | Model Name | Scale | Instruct? | Open-Source? | Checkpoint |
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-All model checkpoints we use and their properties are listed in Tab. 5.
+8B | meta-llama/Meta-Llama-3-8B-Instruct | ||
+8B | meta-llama/Meta-Llama-3.1-8B-Instruct | ||
+7B | Qwen/Qwen2-7B-Instruct | ||
+7B | Qwen/Qwen2.5-7B-Instruct | ||
+12B | mistralai/Mistral-Nemo-Instruct-2407 | ||
+8B | CohereForAI/aya-expanse-8b | ||
+NA | gpt-3.5-turbo-0125 | ||
+NA | gpt-4o-mini-2024-07-18 |
+All model checkpoints we use and their properties are listed in Tab.˜5.
 A.2 Dataset
@@
 | XL-WiC | da, de, en, fr, it, ja, ko, nl, zh | bg, et, fa, hr | pilehvar.github.io/xlwic/ |
-After preprocessing, datapoints for each language split are stored in a single JSON file. Tab. 6 summarizes the supported languages for each dataset and the sources from which they are obtained.
+| XQuAD | ar, de, en, es, ru, tr, vi, zh | el, hi, ro, th | google/xquad |
+After preprocessing, datapoints for each language split are stored in a single JSON file. Tab.˜6 summarizes the supported languages for each dataset and the sources from which they are obtained.
 MGSM
 The original dataset consists of parallel datapoints across all language splits and training/test splits of the same size.
-Example datapoint and the Chat Template for few-shot demonstrations can be found in Fig.
-6.
+Example datapoint and the Chat Template for few-shot demonstrations can be found in Fig.˜6.
 XCOPA
@@
 We then merge both into our XCOPA dataset.
-In XCOPA, there are two types of questions: “cause” and “effect”, each corresponding to a distinct template, as shown in Figs.
-7(a) and 7(b).
+In XCOPA, there are two types of questions: “cause” and “effect”, each corresponding to a distinct template, as shown in Figs.˜7(a) and 7(b).
 XL-WiC
@@
 Among all language splits, Estonian (et) contains the fewest datapoints, with in the training split and in the test split.
-For all other languages, we randomly subsample to match the size of the Estonian split to satisfy the demonstration sampling requirements outlined in Fig.
-2.
+For all other languages, we randomly subsample to match the size of the Estonian split to satisfy the demonstration sampling requirements outlined in Fig.˜2.
 To leverage the attention mechanism of transformers Vaswani et al.
 (2017), we add asterisks around the target word in both sentences to indicate that the LLM needs to disambiguate the meaning of that specific word.
-An example datapoint is shown in Fig.
-8.
+An example datapoint is shown in Fig.˜8.
 COMBINED
@@
 This approach ensures that the total number of datapoints and the overall token count are roughly the same across the original three datasets.
+XQuAD
+The dataset consists of parallel question-answering examples across languages, derived from the same subset of the English SQuAD Rajpurkar et al.
+(2016) v1.1 development set.
+For each language, the context passages, questions, and answers are professionally translated while preserving semantic alignment across languages.
+Each example includes a context paragraph, a question, and a span-based answer.
+An example and the input prompt format used in our experiments are illustrated in Fig.˜9.
 FLORES-101
@@
 CIS are drawn form the filtered FLORES-101 dataset.
-An example is provided in Fig.
-9.
+An example is provided in Fig.˜10.
 A.3 Language
@@
 (2023)—that publicly report the language distribution used during pretraining.
-The top 20 languages and their percentages for each model are listed in Tab.
-7 and Tab.
-8, respectively.
+The top 20 languages and their percentages for each model are listed in Tab.˜7 and Tab.˜8, respectively.
 We take the union of these languages to form what we consider a high-resource language list (in terms of the richness in the LLM pretraining dataset), including the following languages:
@@
 A.3.2 Languages We Evaluate
-Tab.
-9 presents the union of languages supported by the three datasets we evaluated (§ 3, datasets).
+Tab.˜9 presents the union of languages supported by the datasets we evaluated (Tab.˜1).
 Based on whether a language appears in the high-resource language list, we categorized the languages into HRL and LRL groups, with classified as HRL and as LRL.
@@
 | Code | Name in English | HRL/LRL | In Which Dataset(s) | Writing System | Family |
-| bg | Bulgarian | Low | XL-WiC | Cyrillic | Indo-European |
-| bn | Bengali | Low | MGSM | Bengali–Assamese | Indo-European |
-| da | Danish | High | XL-WiC | Latin | Indo-European |
-| de | German | High | MGSM, XL-WiC | Latin | Indo-European |
-| en | English | High | MGSM, XL-WiC, XCOPA | Latin | Indo-European |
-| es | Spanish | High | MGSM | Latin | Indo-European |
-| et | Estonian | Low | XCOPA | Latin | Indo-European |
-| fa | Persian | Low | XL-WiC | Arabic | Indo-European |
-| fr | French | High | MGSM, XL-WiC | Latin | Indo-European |
-| hr | Croatian | Low | XL-WiC | Latin | Indo-European |
-| ht | Haitian | Low | XCOPA | Latin | French Creole |
-| id | Indonesian | High | XCOPA | Latin | Austronesian |
-| it | Italian | High | XL-WiC, XCOPA | Latin | Indo-European |
-| ja | Japanese | High | MGSM, XL-WiC | Kana & Chinese Characters | Japonic |
-| ko | Korean | High | XL-WiC | Hangul | Koreanic |
-| nl | Dutch | High | XL-WiC | Latin | Indo-European |
-| qu | Quechua | Low | XCOPA | Latin | Quechumaran |
-| ru | Russian | High | MGSM | Cyrillic | Indo-European |
-| sw | Swahili | Low | MGSM | Latin | Niger–Congo |
-| ta | Tamil | Low | XCOPA | Tamil | Dravidian |
-| te | Telegu | Low | MGSM | Telegu | Dravidian |
-| th | Thai | Low | MGSM, XCOPA | Thai | Kra–Dai |
-| tr | Turkish | High | XCOPA | Latin | Turkic |
-| vi | Vietnamese | Low | XCOPA | Latin | Austroasiatic |
-| zh | Chinese | High | MGSM, XL-WiC, XCOPA | Chinese Characters | Sino-Tibetan |
+| ar | Arabic | H | XQuAD | Arabic | Afro-Asiatic |
+| bg | Bulgarian | L | XL-WiC | Cyrillic | Indo-European |
+| bn | Bengali | L | MGSM | Bengali–Assamese | Indo-European |
+| da | Danish | H | XL-WiC | Latin | Indo-European |
+| de | German | H | MGSM, XL-WiC, XQuAD | Latin | Indo-European |
+| el | Greek | L | XQuAD | Greek | Indo-European |
+| en | English | H | MGSM, XCOPA, XL-WiC XQuAD | Latin | Indo-European |
+| es | Spanish | H | MGSM, XQuAD | Latin | Indo-European |
+| et | Estonian | L | XCOPA | Latin | Indo-European |
+| fa | Persian | L | XL-WiC | Arabic | Indo-European |
+| fr | French | H | MGSM, XL-WiC | Latin | Indo-European |
+| hi | Hindi | L | XQuAD | Devanagari | Indo-European |
+| hr | Croatian | L | XL-WiC | Latin | Indo-European |
+| ht | Haitian | L | XCOPA | Latin | French Creole |
+| id | Indonesian | H | XCOPA | Latin | Austronesian |
+| it | Italian | H | XL-WiC, XCOPA | Latin | Indo-European |
+| ja | Japanese | H | MGSM, XL-WiC | Kana & Hanzi | Japonic |
+| ko | Korean | H | XL-WiC | Hangul | Koreanic |
+| nl | Dutch | H | XL-WiC | Latin | Indo-European |
+| qu | Quechua | L | XCOPA | Latin | Quechumaran |
+| ro | Romanian | L | XQuAD | Latin | Indo-European |
+| ru | Russian | H | MGSM, XQuAD | Cyrillic | Indo-European |
+| sw | Swahili | L | MGSM | Latin | Niger–Congo |
+| ta | Tamil | L | XCOPA | Tamil | Dravidian |
+| te | Telegu | L | MGSM | Telegu | Dravidian |
+| th | Thai | L | MGSM, XCOPA, XQuAD | Thai | Kra–Dai |
+| tr | Turkish | H | XCOPA, XQuAD | Latin | Turkic |
+| vi | Vietnamese | L | XCOPA, XQuAD | Latin | Austroasiatic |
+| zh | Chinese | H | MGSM, XCOPA, XL-WiC XQuAD | Hanzi | Sino-Tibetan |
 Appendix B Experiment Raw Results
@@
 If a neuron is “activated”, we increment the corresponding counter by , otherwise no action.
-To ensure balanced input across our three datasets, we curated a COMBINED dataset, see § A.2 for details.
+To ensure balanced input across our three datasets, we curated a COMBINED dataset, see §˜A.2 for details.
 After processing the inputs of a single ICL mode, each neuron accumulates an “activated” count.
@@
 We examined the overlaps among the most-activated neurons (specialized neurons) across different ICL modes by calculating the Intersection over Union (IoU) scores.
-For ICL modes and , with specialized neurons denoted as sets and , their overlap is quantified by Eq.
-4:
+For ICL modes and , with specialized neurons denoted as sets and , their overlap is quantified by Eq.˜4:
 Prior research Tang et al.
@@
 In this case, the patterns were less obvious and model-specific.
-This result aligns with our findings in vanilla evaluation that English Non-English HRL Monolingual Multilingual (§ 4.2).
+This result aligns with our findings in vanilla evaluation that English Non-English HRL Monolingual Multilingual (§˜4.2).
 We further split this neuron experiment to either use only LRL or HRL ICL modes when recording neuron activations.
@@
 | Native | 71.54 | 70.26 | 76.15 | 73.33 | 65.13 | 82.82 | 71.79 | 67.95 | 67.95 | 66.15 | 76.41 | 75.64 | 72.05 | |||
+| XQuAD | ar | de | el | en | es | hi | ro | ru | th | tr | vi | zh | LRL Avg | HRL Avg | ALL Avg |
+| English | 60.90 | 75.80 | 69.10 | 86.80 | 79.10 | 69.90 | 79.20 | 62.20 | 69.90 | 72.20 | 73.90 | 77.40 | 72.02 | 73.54 | 73.03 |
+| Chinese | 65.80 | 76.30 | 68.40 | 85.50 | 79.80 | 71.00 | 79.40 | 65.60 | 75.30 | 74.30 | 75.90 | 81.30 | |||
+| German | 63.50 | 75.10 | 69.40 | 87.10 | 79.20 | 70.80 | 79.00 | 62.70 | 72.50 | 73.30 | 76.70 | 77.90 | |||
+| Multilingual | 65.70 | 75.80 | 71.90 | 87.40 | 80.60 | 71.30 | 80.10 | 65.90 | 73.90 | 74.70 | 76.70 | 81.00 | |||
+| Native | 66.70 | 75.10 | 72.00 | 86.80 | 79.40 | 73.60 | 79.90 | 63.90 | 75.40 | 75.60 | 78.60 | 81.30 | |||
+| English | 60.80 | 73.20 | 65.40 | 85.80 | 77.80 | 66.50 | 73.70 | 53.60 | 67.40 | 66.20 | 71.90 | 73.80 | 68.25 | 70.39 | 69.68 |
+| Chinese | 60.00 | 72.50 | 62.70 | 75.70 | 76.20 | 67.60 | 73.20 | 54.50 | 68.40 | 65.30 | 70.70 | 75.00 | |||
+| German | 59.60 | 72.80 | 66.60 | 66.80 | 78.30 | 67.40 | 75.10 | 55.70 | 67.70 | 66.90 | 73.40 | 74.50 | |||
+| Multilingual | 60.80 | 73.50 | 68.10 | 83.50 | 79.70 | 69.40 | 76.00 | 58.20 | 71.00 | 66.60 | 73.00 | 75.10 | |||
+| Native | 61.60 | 72.80 | 67.30 | 85.80 | 77.40 | 69.30 | 76.90 | 59.00 | 71.60 | 67.70 | 75.80 | 75.00 | |||
+| English | 60.40 | 69.60 | 39.10 | 80.70 | 77.20 | 50.80 | 69.60 | 54.90 | 54.10 | 61.00 | 74.40 | 80.10 | 53.40 | 69.79 | 64.32 |
+| Chinese | 58.00 | 68.40 | 36.80 | 83.10 | 74.60 | 48.80 | 67.40 | 53.80 | 56.60 | 61.40 | 72.80 | 83.70 | |||
+| German | 61.00 | 70.40 | 49.20 | 82.40 | 77.10 | 54.10 | 72.30 | 59.60 | 63.10 | 65.80 | 73.60 | 82.20 | |||
+| Multilingual | 60.60 | 71.50 | 48.50 | 83.10 | 78.10 | 55.00 | 72.80 | 60.80 | 64.40 | 64.90 | 74.60 | 82.10 | |||
+| Native | 62.80 | 70.40 | 55.70 | 80.70 | 77.20 | 58.60 | 72.90 | 65.20 | 72.40 | 65.70 | 75.70 | 83.70 | |||
+| English | 63.70 | 74.30 | 57.00 | 85.50 | 80.60 | 62.20 | 80.20 | 67.00 | 76.50 | 69.90 | 75.10 | 85.10 | 68.98 | 75.15 | 73.09 |
+| Chinese | 57.80 | 67.30 | 54.10 | 69.60 | 67.40 | 60.90 | 77.10 | 61.30 | 76.90 | 70.10 | 65.30 | 87.20 | |||
+| German | 63.00 | 75.90 | 57.30 | 82.70 | 80.50 | 62.50 | 79.80 | 67.80 | 76.20 | 71.60 | 76.00 | 85.20 | |||
+| Multilingual | 63.70 | 75.50 | 58.40 | 81.20 | 80.20 | 62.30 | 80.40 | 67.90 | 77.90 | 70.80 | 76.30 | 86.20 | |||
+| Native | 67.90 | 75.90 | 61.00 | 85.50 | 79.80 | 64.90 | 79.50 | 69.30 | 78.40 | 73.10 | 78.10 | 87.20 | |||
+| English | 56.50 | 70.30 | 52.90 | 83.90 | 78.70 | 51.50 | 73.60 | 56.50 | 65.70 | 63.50 | 65.90 | 51.30 | 60.93 | 65.82 | 64.19 |
+| Chinese | 60.40 | 71.80 | 61.30 | 84.30 | 78.40 | 69.90 | 73.70 | 60.80 | 70.60 | 65.60 | 75.50 | 76.2 | |||
+| German | 60.00 | 70.70 | 61.30 | 84.50 | 79.10 | 69.40 | 74.40 | 60.70 | 69.40 | 65.90 | 74.80 | 72.40 | |||
+| Multilingual | 61.30 | 72.30 | 62.60 | 84.80 | 79.60 | 70.00 | 75.40 | 62.40 | 72.30 | 68.00 | 75.20 | 75.20 | |||
+| Native | 61.50 | 70.70 | 63.80 | 83.90 | 76.40 | 70.50 | 76.00 | 63.00 | 72.10 | 68.20 | 74.10 | 76.20 | |||
+| English | 71.60 | 76.70 | 74.00 | 87.60 | 83.70 | 70.30 | 83.00 | 64.70 | 54.60 | 68.40 | 80.20 | 80.60 | 70.48 | 76.69 | 74.62 |
+| Chinese | 72.00 | 72.20 | 72.00 | 66.10 | 82.50 | 69.20 | 81.60 | 63.70 | 54.40 | 63.50 | 79.50 | 82.10 | |||
+| German | 71.60 | 74.70 | 68.40 | 63.40 | 78.20 | 69.50 | 76.70 | 50.60 | 55.10 | 59.50 | 78.20 | 68.70 | |||
+| Multilingual | 72.50 | 74.20 | 72.30 | 78.20 | 83.10 | 71.20 | 82.40 | 64.10 | 55.10 | 67.90 | 79.80 | 78.20 | |||
+| Native | 72.00 | 74.70 | 72.00 | 87.60 | 82.80 | 67.70 | 81.30 | 64.60 | 51.50 | 67.60 | 79.60 | 82.10 | |||
 | McNemar’s Test | Low-Resource Languages | High-Resource Languages | ||||||||||||
-| MGSM ICL Mode | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | ||
-| M1 vs M2 Comparison | Level | Wrong | M2 Correct | M2 Wrong | Correct | Level | False | M2 Correct | M2 Wrong | Correct | ||||
+| MGSM ICL Mode | -value | Sig. | #Both | #M1 | #M1 | #Both | -value | Sig. | #Both | #M1 | #M1 | #Both | ||
+| M1 vs M2 Comparison | Level | M2 | M2 | Level | M2 | M2 | ||||||||
 | English vs French | 0.85 | 280 | 78 | 91 | 551 | 0.08 | 300 | 104 | 109 | 1237 | ||||
@@
 | McNemar’s Test | Low-Resource Languages | High-Resource Languages | ||||||||||||
-| XCOPA ICL Mode | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | ||
-| M1 vs M2 Comparison | Level | Wrong | M2 Correct | M2 Wrong | Correct | Level | Wrong | M2 Correct | M2 Wrong | Correct | ||||
+| XCOPA ICL Mode | -value | Sig. | #Both | #M1 | #M1 | #Both | -value | Sig. | #Both | #M1 | #M1 | #Both | ||
+| M1 vs M2 Comparison | Level | M2 | M2 | Level | M2 | M2 | ||||||||
 | English vs Italian | 274.27 | *** | 1246 | 636 | 166 | 1452 | 8.84 | ** | 287 | 111 | 70 | 2032 | ||
@@
 | McNemar’s Test | Low-Resource Languages | High-Resource Languages | ||||||||||||
-| XL-WiC ICL Mode | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | ||
-| M1 vs M2 Comparison | Level | Wrong | M2 Correct | M2 Wrong | Correct | Level | Wrong | M2 Correct | M2 Wrong | Correct | ||||
+| XL-WiC ICL Mode | -value | Sig. | #Both | #M1 | #M1 | #Both | -value | Sig. | #Both | #M1 | #M1 | #Both | ||
+| M1 vs M2 Comparison | Level | M2 | M2 | Level | M2 | M2 | ||||||||
 | English vs French | 1.85 | 483 | 182 | 156 | 739 | 3.70 | 991 | 422 | 367 | 1730 | ||||
@@
 | English vs Native | 3.20 | 304 | 339 | 293 | 624 | 133.37 | *** | 904 | 676 | 312 | 1618 | |||
+| McNemar’s Test | Low-Resource Languages | High-Resource Languages | ||||||||||||
+| XQuAD ICL Mode | -value | Sig. | #Both | #M1 | #M1 | #Both | -value | Sig. | #Both | #M1 | #M1 | #Both | ||
+| M1 vs M2 Comparison | Level | M2 | M2 | Level | M2 | M2 | ||||||||
+| English vs German | 5.02 | * | 979 | 140 | 104 | 2777 | 10.55 | ** | 1842 | 275 | 203 | 5680 | ||
+| English vs Chinese | 10.61 | ** | 925 | 194 | 134 | 2747 | 42.92 | *** | 1734 | 383 | 221 | 5662 | ||
+| English vs Multilingual | 27.27 | *** | 925 | 194 | 103 | 2778 | 68.30 | *** | 1744 | 373 | 178 | 5705 | ||
+| English vs Native | 42.90 | *** | 867 | 252 | 124 | 2757 | 59.08 | *** | 1716 | 401 | 210 | 5673 | ||
+| English vs German | 4.72 | * | 1106 | 164 | 126 | 2604 | 32.10 | *** | 2094 | 275 | 426 | 5205 | ||
+| English vs Chinese | 0.28 | 1097 | 173 | 184 | 2546 | 25.24 | *** | 2095 | 274 | 406 | 5225 | |||
+| English vs Multilingual | 40.49 | *** | 1052 | 218 | 103 | 2627 | 8.95 | ** | 2043 | 326 | 253 | 5378 | ||
+| English vs Native | 35.73 | *** | 1008 | 262 | 141 | 2589 | 23.68 | *** | 2010 | 359 | 239 | 5392 | ||
+| English vs German | 135.57 | *** | 1508 | 356 | 105 | 2031 | 23.40 | *** | 1947 | 470 | 332 | 5251 | ||
+| English vs Chinese | 3.90 | * | 1689 | 175 | 215 | 1921 | 0.82 | 2077 | 340 | 365 | 5218 | |||
+| English vs Multilingual | 152.19 | *** | 1489 | 375 | 104 | 2032 | 36.32 | *** | 1918 | 499 | 325 | 5258 | ||
+| English vs Native | 275.76 | *** | 1252 | 612 | 152 | 1984 | 52.53 | *** | 1798 | 619 | 388 | 5195 | ||
+| English vs German | 0.00 | 1111 | 130 | 131 | 2628 | 0.39 | 1726 | 262 | 247 | 5765 | ||||
+| English vs Chinese | 13.56 | *** | 1105 | 136 | 205 | 2554 | 288.05 | *** | 1737 | 251 | 803 | 5209 | ||
+| English vs Multilingual | 3.40 | 1093 | 148 | 117 | 2642 | 0.04 | 1695 | 293 | 287 | 5725 | ||||
+| English vs Native | 17.74 | *** | 1030 | 211 | 132 | 2627 | 42.00 | *** | 1624 | 364 | 208 | 5804 | ||
+| English vs German | 177.16 | *** | 1143 | 420 | 112 | 2325 | 182.14 | *** | 2056 | 678 | 263 | 5003 | ||
+| English vs Chinese | 180.26 | *** | 1123 | 440 | 121 | 2316 | 211.57 | *** | 1990 | 744 | 278 | 4988 | ||
+| English vs Multilingual | 217.81 | *** | 1072 | 491 | 124 | 2313 | 286.72 | *** | 1995 | 739 | 215 | 5051 | ||
+| English vs Native | 211.34 | *** | 1017 | 546 | 159 | 2278 | 200.47 | *** | 1939 | 795 | 321 | 4945 | ||
+| English vs German | 47.23 | *** | 1087 | 94 | 216 | 2603 | 459.12 | *** | 1697 | 168 | 854 | 5281 | ||
+| English vs Chinese | 8.30 | ** | 1077 | 104 | 151 | 2668 | 136.10 | *** | 1653 | 212 | 531 | 5604 | ||
+| English vs Multilingual | 0.31 | 1081 | 100 | 109 | 2710 | 42.27 | *** | 1662 | 203 | 358 | 5777 | |||
+| English vs Native | 24.43 | *** | 1051 | 130 | 224 | 2595 | 1.26 | 1649 | 216 | 241 | 5894 | |||
 | MGSM CIS | bn | de | en | es | fr | ja | ru | sw | te | th | zh | LRL AVG | HRL AVG | ALL AVG |
@@
 | English CIS-Multi | 57.95 | 60.77 | 64.36 | 65.64 | 59.49 | 74.36 | 56.41 | 57.18 | 39.49 | 62.31 | 65.64 | 64.87 | 58.21 | 62.24 | 59.74 | 60.51 |
+| XQuAD CIS | ar | de | el | en | es | hi | ro | ru | th | tr | vi | zh | LRL AVG | HRL AVG | ALL AVG |
+| English CIS-En | 60.70 | 75.00 | 68.30 | 86.70 | 79.30 | 68.70 | 79.10 | 62.20 | 69.60 | 72.40 | 74.40 | 77.80 | 71.43 | 73.56 | 72.85 |
+| English CIS-Fr | 58.80 | 75.30 | 68.30 | 86.70 | 79.50 | 68.00 | 79.00 | 62.00 | 68.60 | 71.30 | 73.60 | 76.70 | 70.98 | 72.99 | 72.32 |
+| English CIS-Ja | 58.60 | 75.00 | 67.70 | 86.80 | 79.70 | 66.90 | 79.20 | 61.70 | 69.60 | 71.50 | 73.00 | 77.80 | 70.85 | 73.01 | 72.29 |
+| English CIS-Zh | 58.30 | 75.30 | 68.40 | 87.00 | 79.40 | 68.20 | 78.80 | 61.90 | 69.70 | 71.40 | 73.20 | 77.90 | 71.28 | 73.05 | 72.46 |
+| English CIS-Multi | 58.50 | 75.30 | 68.70 | 87.30 | 79.80 | 66.40 | 79.20 | 62.30 | 69.30 | 71.60 | 74.10 | 76.70 | 70.90 | 73.20 | 72.43 |
+| English CIS-En | 60.30 | 73.20 | 65.50 | 84.40 | 78.20 | 67.10 | 74.60 | 54.70 | 66.60 | 67.10 | 70.90 | 74.20 | 68.45 | 70.38 | 69.73 |
+| English CIS-Fr | 61.00 | 73.10 | 65.30 | 85.20 | 77.80 | 67.40 | 74.50 | 55.70 | 68.60 | 65.80 | 71.90 | 73.40 | 68.95 | 70.49 | 69.98 |
+| English CIS-Ja | 60.80 | 72.60 | 65.00 | 84.30 | 77.30 | 67.80 | 74.80 | 55.30 | 68.00 | 66.70 | 70.70 | 74.70 | 68.90 | 70.30 | 69.83 |
+| English CIS-Zh | 61.10 | 73.00 | 65.90 | 84.70 | 77.30 | 67.70 | 74.60 | 55.80 | 68.20 | 66.20 | 71.30 | 74.40 | 69.10 | 70.48 | 70.02 |
+| English CIS-Multi | 61.00 | 72.80 | 66.30 | 84.80 | 77.90 | 67.90 | 75.00 | 55.60 | 69.20 | 66.10 | 71.40 | 73.20 | 69.60 | 70.35 | 70.10 |
+| English CIS-En | 61.00 | 70.50 | 47.70 | 80.90 | 77.20 | 54.50 | 72.30 | 57.30 | 59.70 | 63.00 | 73.70 | 81.80 | 58.55 | 70.68 | 66.63 |
+| English CIS-Fr | 60.30 | 71.10 | 50.80 | 81.40 | 77.00 | 55.70 | 73.00 | 59.50 | 64.10 | 64.10 | 73.10 | 81.80 | 60.90 | 71.04 | 67.66 |
+| English CIS-Ja | 60.80 | 69.70 | 50.10 | 82.40 | 77.50 | 55.10 | 72.10 | 58.90 | 65.00 | 63.80 | 74.10 | 81.50 | 60.58 | 71.09 | 67.58 |
+| English CIS-Zh | 60.40 | 69.40 | 47.80 | 81.20 | 75.90 | 54.30 | 71.50 | 57.90 | 62.20 | 62.40 | 72.90 | 81.70 | 58.95 | 70.22 | 66.47 |
+| English CIS-Multi | 59.80 | 70.10 | 50.00 | 82.00 | 77.80 | 54.40 | 72.30 | 59.80 | 64.60 | 64.50 | 73.50 | 81.40 | 60.32 | 71.11 | 67.52 |
+| English CIS-En | 64.20 | 75.60 | 56.90 | 85.60 | 80.50 | 62.60 | 80.30 | 66.20 | 76.40 | 70.60 | 75.40 | 85.20 | 69.05 | 75.41 | 73.29 |
+| English CIS-Fr | 62.80 | 74.90 | 57.30 | 86.10 | 80.60 | 62.60 | 80.20 | 66.40 | 76.40 | 70.10 | 75.60 | 85.10 | 69.12 | 75.20 | 73.17 |
+| English CIS-Ja | 58.40 | 73.40 | 57.70 | 85.10 | 77.60 | 62.20 | 79.90 | 64.10 | 74.80 | 69.80 | 71.50 | 85.70 | 68.65 | 73.20 | 71.68 |
+| English CIS-Zh | 60.40 | 73.90 | 56.50 | 85.40 | 77.80 | 61.90 | 78.80 | 64.30 | 75.80 | 69.70 | 70.50 | 85.70 | 68.25 | 73.46 | 71.72 |
+| English CIS-Multi | 62.80 | 74.60 | 56.90 | 85.90 | 80.10 | 61.80 | 79.80 | 66.50 | 76.00 | 69.40 | 74.90 | 85.50 | 68.62 | 74.96 | 72.85 |
+| English CIS-En | 54.40 | 69.60 | 46.70 | 82.90 | 76.70 | 44.30 | 71.40 | 55.20 | 63.50 | 61.70 | 64.10 | 49.50 | 56.48 | 64.26 | 61.67 |
+| English CIS-Fr | 47.80 | 63.00 | 41.90 | 83.50 | 75.50 | 37.20 | 66.30 | 43.40 | 59.20 | 53.90 | 53.40 | 37.20 | 51.15 | 57.21 | 55.19 |
+| English CIS-Ja | 46.20 | 63.30 | 40.70 | 83.00 | 75.60 | 33.40 | 66.10 | 45.40 | 59.40 | 55.20 | 54.60 | 34.90 | 49.90 | 57.28 | 54.82 |
+| English CIS-Zh | 46.30 | 62.70 | 42.50 | 83.10 | 75.10 | 33.00 | 66.50 | 41.70 | 58.20 | 55.10 | 54.40 | 32.70 | 50.05 | 56.39 | 54.28 |
+| English CIS-Multi | 47.00 | 63.50 | 42.00 | 83.30 | 75.00 | 37.60 | 67.90 | 43.70 | 59.90 | 56.00 | 54.10 | 39.90 | 51.85 | 57.81 | 55.82 |
+| English CIS-En | 72.00 | 77.20 | 73.80 | 88.30 | 84.10 | 70.40 | 82.10 | 64.30 | 54.90 | 68.50 | 79.70 | 81.40 | 70.30 | 76.94 | 74.73 |
+| English CIS-Fr | 72.10 | 75.60 | 72.40 | 88.50 | 84.30 | 70.30 | 81.30 | 63.70 | 54.70 | 66.80 | 78.60 | 78.60 | 69.68 | 76.02 | 73.91 |
+| English CIS-Ja | 71.40 | 75.90 | 72.70 | 88.20 | 83.80 | 69.90 | 82.20 | 64.40 | 55.20 | 68.80 | 78.50 | 77.30 | 70.00 | 76.04 | 74.02 |
+| English CIS-Zh | 71.10 | 76.40 | 73.10 | 88.70 | 83.90 | 70.40 | 82.50 | 64.80 | 55.00 | 68.70 | 79.00 | 81.70 | 70.25 | 76.79 | 74.61 |
+| English CIS-Multi | 71.70 | 76.80 | 73.00 | 88.50 | 83.80 | 70.60 | 82.30 | 64.00 | 54.50 | 67.70 | 79.60 | 80.40 | 70.10 | 76.56 | 74.41 |
 | McNemar’s Test | Low-Resource Languages | High-Resource Languages | ||||||||||||
-| MGSM CIS Mode | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | ||
-| M1 vs M2 Comparison | Level | Wrong | M2 Correct | M2 Wrong | Correct | Level | False | M2 Correct | M2 Wrong | Correct | ||||
+| MGSM CIS Mode | -value | Sig. | #Both | #M1 | #M1 | #Both | -value | Sig. | #Both | #M1 | #M1 | #Both | ||
+| M1 vs M2 Comparison | Level | M2 | M2 | Level | M2 | M2 | ||||||||
 | English CIS-En vs English CIS-Zh | 1.57 | 2.10E-01 | 317 | 73 | 90 | 520 | 0.47 | 4.93E-01 | 339 | 91 | 81 | 1239 | ||
@@
 | McNemar’s Test | Low-Resource Languages | High-Resource Languages | ||||||||||||
-| XCOPA CIS Mode | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | ||
-| M1 vs M2 Comparison | Level | Wrong | M2 Correct | M2 Wrong | Correct | Level | False | M2 Correct | M2 Wrong | Correct | ||||
+| XCOPA CIS Mode | -value | Sig. | #Both | #M1 | #M1 | #Both | -value | Sig. | #Both | #M1 | #M1 | #Both | ||
+| M1 vs M2 Comparison | Level | M2 | M2 | Level | M2 | M2 | ||||||||
 | English CIS-En vs English CIS-Zh | 199.24 | 3.05E-45 | *** | 1353 | 454 | 116 | 1577 | 8.28 | 4.00E-03 | ** | 349 | 58 | 30 | 2063 |
@@
 | McNemar’s Test | Low-Resource Languages | High-Resource Languages | ||||||||||||
-| XL-WiC CIS Mode | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | -value | Sig. | #Both | #M1 Wrong | #M1 Correct | #Both | ||
-| M1 vs M2 Comparison | Level | Wrong | M2 Correct | M2 Wrong | Correct | Level | False | M2 Correct | M2 Wrong | Correct | ||||
+| XL-WiC CIS Mode | -value | Sig. | #Both | #M1 | #M1 | #Both | -value | Sig. | #Both | #M1 | #M1 | #Both | ||
+| M1 vs M2 Comparison | Level | M2 | M2 | Level | M2 | M2 | ||||||||
 | English CIS-En vs English CIS-Zh | 3.31 | 6.90E-02 | 583 | 92 | 68 | 817 | 1.69 | 1.94E-01 | 1251 | 183 | 158 | 1918 | ||
@@
 | Multilingual CIS-Multi vs English CIS-Multi | 0.80 | 3.72E-01 | 417 | 190 | 172 | 781 | 43.63 | 3.96E-11 | *** | 965 | 270 | 448 | 1827 | |
+| McNemar’s Test | Low-Resource Languages | High-Resource Languages | ||||||||||||
+| XQuAD CIS Mode | -value | Sig. | #Both | #M1 | #M1 | #Both | -value | Sig. | #Both | #M1 | #M1 | #Both | ||
+| M1 vs M2 Comparison | Level | M2 | M2 | Level | M2 | M2 | ||||||||
+| English CIS-En vs English CIS-Zh | 0.19 | 6.66E-01 | 1079 | 64 | 70 | 2787 | 6.53 | 1.06E-02 | * | 2013 | 102 | 143 | 5742 | |
+| English CIS-En vs English CIS-Fr | 2.37 | 1.24E-01 | 1091 | 52 | 70 | 2787 | 8.80 | 3.01E-03 | ** | 2023 | 92 | 138 | 5747 | |
+| English CIS-En vs English CIS-Ja | 3.21 | 7.34E-02 | 1079 | 64 | 87 | 2770 | 7.06 | 7.89E-03 | ** | 2006 | 109 | 153 | 5732 | |
+| English CIS-En vs English CIS-Multi | 2.55 | 1.10E-01 | 1075 | 68 | 89 | 2768 | 3.23 | 7.25E-02 | 2008 | 107 | 136 | 5749 | ||
+| Multilingual CIS-Multi vs English CIS-Multi | 51.05 | 9.01E-13 | *** | 945 | 92 | 219 | 2744 | 65.38 | 6.18E-16 | *** | 1769 | 183 | 375 | 5673 |
+| English CIS-En vs English CIS-Zh | 3.47 | 6.24E-02 | 1159 | 103 | 77 | 2661 | 0.15 | 6.98E-01 | 2203 | 167 | 159 | 5471 | ||
+| English CIS-En vs English CIS-Fr | 2.44 | 1.18E-01 | 1178 | 84 | 64 | 2674 | 0.25 | 6.20E-01 | 2235 | 135 | 126 | 5504 | ||
+| English CIS-En vs English CIS-Ja | 1.54 | 2.15E-01 | 1159 | 103 | 85 | 2653 | 0.07 | 7.86E-01 | 2204 | 166 | 172 | 5458 | ||
+| English CIS-En vs English CIS-Multi | 11.13 | 8.51E-04 | *** | 1148 | 114 | 68 | 2670 | 0.00 | 9.55E-01 | 2216 | 154 | 156 | 5474 | |
+| Multilingual CIS-Multi vs English CIS-Multi | 15.06 | 1.04E-04 | *** | 1066 | 89 | 150 | 2695 | 18.72 | 1.51E-05 | *** | 2083 | 193 | 289 | 5435 |
+| English CIS-En vs English CIS-Zh | 1.14 | 2.86E-01 | 1551 | 107 | 91 | 2251 | 3.83 | 5.04E-02 | 2204 | 142 | 178 | 5476 | ||
+| English CIS-En vs English CIS-Fr | 40.04 | 2.49E-10 | *** | 1503 | 155 | 61 | 2281 | 2.37 | 1.24E-01 | 2166 | 180 | 151 | 5503 | |
+| English CIS-En vs English CIS-Ja | 25.50 | 4.43E-07 | *** | 1492 | 166 | 85 | 2257 | 2.76 | 9.66E-02 | 2144 | 202 | 169 | 5485 | |
+| English CIS-En vs English CIS-Multi | 23.90 | 1.01E-06 | *** | 1520 | 138 | 67 | 2275 | 3.31 | 6.88E-02 | 2154 | 192 | 157 | 5497 | |
+| Multilingual CIS-Multi vs English CIS-Multi | 13.06 | 3.02E-04 | *** | 1403 | 120 | 184 | 2293 | 8.32 | 3.91E-03 | ** | 1990 | 251 | 321 | 5438 |
+| English CIS-En vs English CIS-Zh | 6.08 | 1.37E-02 | * | 1175 | 63 | 95 | 2667 | 54.60 | 1.48E-13 | *** | 1825 | 142 | 298 | 5735 |
+| English CIS-En vs English CIS-Fr | 0.03 | 8.55E-01 | 1177 | 61 | 58 | 2704 | 1.00 | 3.16E-01 | 1848 | 119 | 136 | 5897 | ||
+| English CIS-En vs English CIS-Ja | 1.29 | 2.55E-01 | 1159 | 79 | 95 | 2667 | 63.09 | 1.98E-15 | *** | 1810 | 157 | 334 | 5699 | |
+| English CIS-En vs English CIS-Multi | 1.77 | 1.84E-01 | 1174 | 64 | 81 | 2681 | 4.06 | 4.40E-02 | * | 1834 | 133 | 169 | 5864 | |
+| Multilingual CIS-Multi vs English CIS-Multi | 6.34 | 1.18E-02 | * | 1115 | 100 | 140 | 2645 | 0.05 | 8.31E-01 | 1724 | 273 | 279 | 5724 | |
+| English CIS-En vs English CIS-Zh | 175.70 | 4.21E-40 | *** | 1683 | 58 | 315 | 1944 | 462.20 | 1.60E-102 | *** | 2746 | 113 | 743 | 4398 |
+| English CIS-En vs English CIS-Fr | 125.89 | 3.24E-29 | *** | 1669 | 72 | 285 | 1974 | 380.06 | 1.21E-84 | *** | 2724 | 135 | 699 | 4442 |
+| English CIS-En vs English CIS-Ja | 175.56 | 4.52E-40 | *** | 1677 | 64 | 327 | 1932 | 380.18 | 1.14E-84 | *** | 2729 | 130 | 689 | 4452 |
+| English CIS-En vs English CIS-Multi | 98.13 | 3.91E-23 | *** | 1661 | 80 | 265 | 1994 | 329.88 | 1.02E-73 | *** | 2715 | 144 | 660 | 4481 |
+| Multilingual CIS-Multi vs English CIS-Multi | 562.57 | 2.32E-124 | *** | 1140 | 85 | 786 | 1989 | 871.12 | 1.86E-191 | *** | 2073 | 169 | 1302 | 4456 |
+| English CIS-En vs English CIS-Zh | 0.01 | 9.17E-01 | 1143 | 45 | 47 | 2765 | 0.78 | 3.78E-01 | 1773 | 72 | 84 | 6071 | ||
+| English CIS-En vs English CIS-Fr | 5.49 | 1.92E-02 | * | 1148 | 40 | 65 | 2747 | 20.99 | 4.62E-06 | *** | 1758 | 87 | 160 | 5995 |
+| English CIS-En vs English CIS-Ja | 1.12 | 2.90E-01 | 1140 | 48 | 60 | 2752 | 18.81 | 1.44E-05 | *** | 1747 | 98 | 170 | 5985 | |
+| English CIS-En vs English CIS-Multi | 0.57 | 4.50E-01 | 1149 | 39 | 47 | 2765 | 6.01 | 1.42E-02 | * | 1790 | 55 | 85 | 6070 | |
+| Multilingual CIS-Multi vs English CIS-Multi | 1.28 | 2.58E-01 | 1116 | 96 | 80 | 2708 | 39.10 | 4.02E-10 | *** | 1704 | 309 | 171 | 5816 | |